### v2

# Tight Guarantees for Cut-Relative Survivable Network Design via a Decomposition Technique

Abstract
In the classical survivable-network-design problem ($\mathsf{SNDP}$), we are given an undirected graph $G=(V,E)$, non-negative edge costs, and some $k$ tuples $(s_{i},t_{i},r_{i})$, where $s_{i},t_{i}\in V$ and $r_{i}\in\mathbb{Z}_{+}$. The objective is to find a minimum-cost subset $H\subseteq E$ such that each $s_{i}$-$t_{i}$ pair remains connected even after the failure of any $r_{i}-1$ edges. It is well-known that $\mathsf{SNDP}$ can be equivalently modeled using a weakly-supermodular cut-requirement function $f$, where the objective is to find the minimum-cost subset of edges that picks at least $f(S)$ edges across every cut $S\subseteq V$.
Recently, motivated by fault-tolerance in graph spanners, Dinitz, Koranteng, and Kortsartz proposed a variant of $\mathsf{SNDP}$ that enforces a relative level of fault tolerance with respect to $G$. Even if a feasible $\mathsf{SNDP}$-solution may not exist due to $G$ lacking the required fault-tolerance, the goal is to find a solution $H$ that is at least as fault-tolerant as $G$ itself. They formalize the latter condition in terms of paths and fault-sets, which gives rise to path-relative $\mathsf{SNDP}$ (which they call relative $\mathsf{SNDP}$). Along these lines, we introduce a new model of relative network design, called cut-relative $\mathsf{SNDP}$ ($\mathsf{CR}\text{-}\mathsf{SNDP}$), where the goal is to select a minimum-cost subset of edges that satisfies the given (weakly-supermodular) cut-requirement function to the maximum extent possible, i.e., by picking $\min\{f(S),|\delta_{G}(S)|\}$ edges across every cut $S\subseteq V$.
Unlike $\mathsf{SNDP}$, the cut-relative and path-relative versions of $\mathsf{SNDP}$ are not equivalent. The resulting cut-requirement function for $\mathsf{CR}\text{-}\mathsf{SNDP}$ (as also path-relative $\mathsf{SNDP}$) is not weakly supermodular, and extreme-point solutions to the natural LP-relaxation need not correspond to a laminar family of tight cut constraints. Consequently, standard techniques cannot be used directly to design approximation algorithms for this problem. We develop a novel decomposition technique to circumvent this difficulty and use it to give a tight $2$-approximation algorithm for $\mathsf{CR}\text{-}\mathsf{SNDP}$. We also show some new hardness results for these relative-$\mathsf{SNDP}$ problems.
keywords:
Approximation algorithms, Network Design, Cut-requirement functions, Weak Supermodularity, Iterative rounding, LP rounding algorithms
1 Introduction
In the classical survivable-network-design problem ($\mathsf{SNDP}$), we are given an undirected graph $G=(V,E)$, non-negative edge costs $\{c_{e}\}_{e\in E}$, and some $k$ source-sink pairs $s_{i},t_{i}\in V$ with requirements $r_{i}\in\mathbb{Z}_{+}$, for $i=1,\ldots,k$. The goal is to find a minimum-cost set $H\subseteq E$ of edges such that there are at least $r_{i}$ edge-disjoint $s_{i}$-$t_{i}$ paths in $H$, for all $i\in[k]$. We overload notation and use $H$ to denote both the edge-set, and the corresponding subgraph of $G$.
$\mathsf{SNDP}$ imposes an absolute level of fault-tolerance in a solution subgraph $H$, by ensuring that for any bounded-size fault-set $F\subseteq E$ of edges that may fail, the graph $H-F$ retains some connectivity properties. More precisely, due to Menger’s theorem, or the max-flow min-cut theorem, the feasibility condition on $H$ can be stated equivalently as follows. For every $i\in[k]$, and every fault-set $F\subseteq E$ with $|F|<r_{i}$:
  1. (S1)
    (Path-version) there is an $s_{i}$-$t_{i}$ path in $H-F$;
  2. (S2)
    (Cut-version) for every $s_{i}$-$t_{i}$ cut $S\subseteq V$ (i.e., $|S\cap\{s_{i},t_{i}\}|=1$), we have $\delta_{H-F}(S)\neq\emptyset$, where $\delta_{H-F}(S):=\delta(S)\cap(H-F)$.
Recently, motivated by work on fault-tolerance in graph spanners, Dinitz, Koranteng, and Kortsartz [8], proposed a variant of $\mathsf{SNDP}$ that aims to impose a relative level of fault-tolerance with respect to the graph $G$. The idea is that even if the $\mathsf{SNDP}$ instance in infeasible, because $G$ itself does not possess the required level of fault-tolerance, one should not have to completely abandon the goal of fault-tolerance: it is still meaningful and reasonable to seek a solution that is “as fault-tolerant as $G$”, and in this sense is fault-tolerant relative to $G$.
To formalize this, it is useful to consider the definition of $\mathsf{SNDP}$ in terms of fault-sets. Roughly speaking, we would like to capture that $H$ is a feasible solution if
for every fault-set $F$ (valid for $\mathsf{SNDP}$),  $H-F$ and $G-F$ have similar connectivity. (*)
To elaborate, in $\mathsf{SNDP}$, if there is even one fault-set $F^{\prime}$ under which $G-F^{\prime}$ fails to have the required connectivity, i.e., $|F^{\prime}|<r_{i}$ and $F^{\prime}\supseteq\delta_{G}(S)$ for some $s_{i}$-$t_{i}$ cut $S$, then “all bets are off”; we declare that the instance in infeasible. But declaring infeasibility here feels unsatisfactory because we allow one specific “problematic” fault-set $F^{\prime}$ to block us from obtaining any kind of fault-tolerance. In contrast, (*1) aims to provide a per-fault-set guarantee, by asking for a solution $H$ that functions as well as $G$ in terms of connectivity under the failure of any fault-set $F$ (considered for $\mathsf{SNDP}$). We can formalize “$H-F$ and $G-F$ have similar connectivity” in two ways, via paths or cuts, and this gives rise to the following two problem definitions.
Definition 1.1.
Let $\bigl{(}G=(V,E),\{c_{e}\}_{e\in E},\{s_{i},t_{i},r_{i}\}_{i\in[k]}\bigr{)}$ be an $\mathsf{SNDP}$ instance.
  1. (R1)
    Path-relative $\mathsf{SNDP}$: $H\subseteq E$ is a feasible solution if for every $i\in[k]$ and $F\subseteq E$ with $|F|<r_{i}$,   $G-F$ has an $s_{i}$-$t_{i}$ path   $\implies$   $H-F$ has an $s_{i}$-$t_{i}$ path.
  2. (R2)
    Cut-relative $\mathsf{SNDP}$ ($\mathsf{CR}\text{-}\mathsf{SNDP}$): $H\subseteq E$ is a feasible solution if for every $i\in[k]$ and $F\subseteq E$ with $|F|<r_{i}$, for every $s_{i}$-$t_{i}$ cut $S$,  $\delta_{G-F}(S)\neq\emptyset\ \implies\ \delta_{H-F}(S)\neq\emptyset$.
The goal in both problems is to find a minimum-cost feasible solution.
Perhaps surprisingly, path-relative $\mathsf{SNDP}$ and cut-relative $\mathsf{SNDP}$ are not equivalent, even when the $\mathsf{SNDP}$ instance involves a single $s$-$t$ pair. It is not hard to show that if $H$ is feasible for $\mathsf{CR}\text{-}\mathsf{SNDP}$, it is also feasible for path-relative SNDP [^1] Fix any $i\in[k]$ and $F\subseteq E$ with $|F|<r_{i}$. If $G-F$ has an $s_{i}$-$t_{i}$ path, then $\delta_{G-F}(S)\neq\emptyset$ for every $s_{i}$-$t_{i}$ cut. So since $H$ is feasible for cut-relative $\mathsf{SNDP}$, we have $\delta_{H-F}(S)\neq\emptyset$ for every $s_{i}$-$t_{i}$ cut, which implies that $H-F$ has an $s_{i}$-$t_{i}$ path. but the converse fails to hold; see Appendix A for an example.
Path-relative $\mathsf{SNDP}$ was defined by Dinitz et al. [8], who referred to this problem simply as relative $\mathsf{SNDP}$. They considered (among other problems) the path-relative version of $k$-edge connected subgraph ($k\text{-}\mathsf{ECSS}$), which is the special case of $\mathsf{SNDP}$ where we have an $s_{i}$-$t_{i}$ pair with $r_{i}=k$ for every pair of nodes. (In this case, the path-relative and cut-relative versions do turn out to be equivalent [8]; see also Lemma A.1.) They called the resulting path-relative problem, $k$-edge fault-tolerant subgraph ($k\text{-}\mathsf{EFTS}$), and observed that the feasibility condition for $k\text{-}\mathsf{EFTS}$ can be equivalently stated as: $H$ is feasible iff $|\delta_{H}(S)|\geq g(S):=\min\bigl{\{}k,|\delta_{G}(S)|\bigr{\}}$ for all $S\subseteq V$. The function $g:2^{V}\mapsto\mathbb{Z}$ is called a cut-requirement function, as it stipulates the (minimum) number of edges across any cut in any feasible solution. Cut-requirement functions constitute a very versatile framework for specifying network-design problems, where given a cut-requirement function $f:2^{V}\mapsto\mathbb{Z}$, the corresponding $f$-network-design problem (${f}\text{-}\mathsf{NDP}$), also called the $f$-connectivity problem, is to find a minimum-cost set $H\subseteq E$ such that $|\delta_{H}(S)|\geq f(S)$ for all $S\subseteq V$. For instance, it is easy to see that $\mathsf{SNDP}$ corresponds to ${f}\text{-}\mathsf{NDP}$ for the cut-requirement function $f^{\mathsf{SNDP}}$, defined by $f^{\mathsf{SNDP}}(S):=\max\,\{r_{i}:|S\cap\{s_{i},t_{i}\}|=1\}$ (and $k\text{-}\mathsf{ECSS}$ is ${f^{k\text{-}\mathsf{ECSS}}}\text{-}\mathsf{NDP}$ where $f^{k\text{-}\mathsf{ECSS}}(S):=k$ for all $\emptyset\neq S\subsetneq V$); various other network-design problems, such as the $T$-join problem, point-to-point connection problem etc., can be modeled using suitable cut-requirement functions (see, e.g., [10, 11]).
Cut-relative $\mathsf{SNDP}$ is a problem that we introduce in this paper. Its formulation in terms of cuts is the problem one obtains when we replace $f^{k\text{-}\mathsf{ECSS}}$ in the cut-based formulation of $k\text{-}\mathsf{EFTS}$ by the cut-requirement function $f^{\mathsf{SNDP}}$ for (general) $\mathsf{SNDP}$: that is, (similar to $k\text{-}\mathsf{EFTS}$), we can say that $H\subseteq E$ is feasible for $\mathsf{CR}\text{-}\mathsf{SNDP}$ iff $|\delta_{H}(S)|\geq g^{\mathsf{CR}\text{-}\mathsf{SNDP}}(S):=\min\bigl{\{}f^{\mathsf{SNDP}}(S),|\delta_{G}(S)|\bigr{\}}$ for all $S\subseteq V$ (we show this easy equivalence in Section 2); in other words, $\mathsf{CR}\text{-}\mathsf{SNDP}$ corresponds to ${g^{\mathsf{CR}\text{-}\mathsf{SNDP}}}\text{-}\mathsf{NDP}$. To gain some intuition for cut-relative $\mathsf{SNDP}$, and contrast it with path-relative $\mathsf{SNDP}$, consider $\mathsf{SNDP}$ with a single $s$-$t$ pair and requirement $r$, for simplicity. A feasible solution $H$ to path-relative $\mathsf{SNDP}$ offers a per-fault-set guarantee as encapsulated by (R1). But, for a given fault-set $F$ with $|F|<r$, we get a weak fault-tolerance guarantee in terms of $s$-$t$ cuts: if $\delta_{G-F}(S)=\emptyset$ for even one $s$-$t$ cut $S$ then there is no requirement on $H$ for this fault-set $F$. Cut-relative $\mathsf{SNDP}$ offers a per-fault-set and per-cut guarantee, since for (every $F$ with $|F|<r$ and) every $s$-$t$ cut $S$, we require that $\delta_{H-F}(S)\neq\emptyset$ if $\delta_{G-F}(S)\neq\emptyset$.
Along the lines of $\mathsf{CR}\text{-}\mathsf{SNDP}$, we can easily extend the cut-relative model to capture, more broadly, the cut-relative version of any network-design problem specified by a cut-requirement function: given a cut-requirement function $f$, in the corresponding cut-relative $f$-network-design problem ($\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$), we seek a min-cost $H\subseteq E$ such that $|\delta_{H}(S)|\geq g^{\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}}(S):=\min\bigl{\{}f(S),|\delta_{G}(S)|\bigr{\}}$ for all $S\subseteq V$; that is, we seek to satisfy the cut-requirement function $f$ to the maximum extent possible. We refer to $f$ as the base cut-requirement function, to distinguish it from the cut-requirement function $g^{\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}}$ that defines $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$.
Modeling power.
We arrived at $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$, which can be seen as a fault-tolerant cut-covering problem, as a technically natural extension of $k\text{-}\mathsf{EFTS}$. We show below that $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ offers a surprising amount of modeling power, and, in particular, allows one to capture stronger forms of relative fault-tolerance compared to $k\text{-}\mathsf{EFTS}$. In $k\text{-}\mathsf{EFTS}$, there is a sharp relative-fault-tolerance threshold at $k$: if $H$ is feasible, then $G-F$ and $H-F$ have the same components whenever $|F|<k$, but there are no guarantees for larger fault-sets. With $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$, one can capture a weaker guarantee also for other fault-sets, which allows for a more graceful degradation of relative fault-tolerance as $|F|$ increases.
Example 1.2.
One way of modeling graceful relative-fault-tolerance degradation is as follows: given a non-increasing function $\tau:\mathbb{Z}_{+}\mapsto\mathbb{Z}_{+}$, we seek a min-cost subgraph $H$ such that for every fault-set $F$, $G-F$ and $H-F$ have exactly the same components with at most $\tau(|F|)$ nodes. If $G$ models a communication network where connected nodes can communicate with each other, then this yields the desirable guarantee that, post-faults, if a node can communicate with more than $\tau(|F|)$ nodes in $G$ then the same holds for $H$; if this number is at most $\tau(|F|)$, then post-faults, it can communicate with the same nodes in $G$ and $H$. Note that if $\tau(k-1)\geq n=|V|$, then $H$ is feasible for $k\text{-}\mathsf{EFTS}$; by adding more “levels” to $\tau$, we can obtain fault-tolerance guarantees for larger fault-sets.
In Section 3, we show that this problem can be modeled as $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ with the weakly-supermodular function $f=f^{\mathsf{grace}}$, where $f^{\mathsf{grace}}(S):=\min\,\{\ell:\tau(\ell)<|S|\}$, for $S\neq\emptyset$.
Example 1.3.
Generalizing Example 1.2 substantially, suppose along with $\tau$, we have a monotone function , i.e., if $T\subseteq S$. We now seek a subgraph $H$ such that for every fault-set $F$ and $S\subseteq V$ with , we have that $S$ is (the node-set of) a component of $G-F$ iff it is a component of $H-F$. Similar to Example 1.2, we can model this as $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ by defining $f(S):=\min\,\{\ell:\tau(\ell)<$, for $S\neq\emptyset$ (Theorem 3.3).
This setup yields fault-tolerance degradation depending on monotone properties of components other than just their size, which creates a rich space of problems. For example, suppose is the weak-diameter of $S$. Setting $\tau(\ell)=n$ if $\ell<k$, and some $t<n$ for larger $\ell$, we seek a $k\text{-}\mathsf{EFTS}$ solution $H$ such that for $F\subseteq E$ with $|F|\geq k$, $G-F$ and $H-F$ have the same components of weak-diameter at most $t$.
1.1 Our contributions
We introduce cut-relative network-design problems and develop strong approximation guarantees for these problems. We obtain an approximation guarantee of $2$ for cut-relative $\mathsf{SNDP}$, which, notably, matches the best-known approximation factor for $\mathsf{SNDP}$. Our guarantee applies more generally to $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$, whenever the base cut-requirement function $f$ satisfies certain properties and the natural LP-relaxation of the problem $(\text{{CRLP}}_{f,G})$ can be solved efficiently. Our guarantee is relative to the optimal value of this LP, and is tight in that it matches the integrality gap of this LP.
We also show that even in the simplest $\mathsf{SNDP}$ setting with only one $s$-$t$ pair (wherein $\mathsf{SNDP}$ is polytime solvable), cut-relative $\mathsf{SNDP}$ and path-relative $\mathsf{SNDP}$ capture $k\text{-}\mathsf{ECSS}$ as a special case, and are thus APX-hard (Section 7). Previously, even NP-hardness of path-relative $\mathsf{SNDP}$ in the $s$-$t$ case (which is studied by [8, 9]) was not known.
Technical contributions and overview.
Technically, our main contribution is to show that we can obtain such a strong guarantee despite the fact that the cut-requirement function defining cut-relative network-design (i.e., $g^{\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}}$) is not weakly supermodular, which is the key property that drives the $2$-approximation algorithm for $\mathsf{SNDP}$. To elaborate, there is a celebrated $2$-approximation algorithm for $\mathsf{SNDP}$ due to Jain [15] based on iterative rounding. This crucially exploits the fact that the underlying cut-requirement function $f^{\mathsf{SNDP}}$ is weakly supermodular: for any two node-sets $A,B\subseteq V$, we have:
$f^{\mathsf{SNDP}}(A)+f^{\mathsf{SNDP}}(B)\leq\max\bigl{\{}f^{\mathsf{SNDP}}(A\cap B)+f^{\mathsf{SNDP}}(A\cup B),\ f^{\mathsf{SNDP}}(A-B)+f^{\mathsf{SNDP}}(B-A)\}.$
This property allows one to argue that given an LP-solution $x$, any two tight sets that cross—i.e., sets $A,B$ for which $x(S)=f^{\mathsf{SNDP}}(S)$ holds for $S\in\{A,B\}$ and $A\cap B,A-B,B-A$ are all non-empty—can be “uncrossed”, and thereby an extreme-point solution $\widehat{x}$ to the LP can be defined via a laminar family of tight cut constraints. By a laminar family, we mean that any two sets $A,B$ in the family satisfy $A\cap B=\emptyset$ or $A\subseteq B$, and uncrossing two crossing tight sets means that they can be replaced by an equivalent laminar family of tight sets. Jain’s seminal contribution was to show that given such a laminar family defining an extreme point, there always exists an edge $e$ for which $\widehat{x}_{e}\geq\frac{1}{2}$; picking such an edge and iterating then yields the $2$-approximation. This technique of uncrossing tight cut-constraints to obtain a laminar family of tight cut-constraints has proved to be quite versatile, becoming a staple technique that has been leveraged to obtain strong guarantees in various network-design settings such as, most notably, network-design problems with degree constraints [13, 17, 19, 5, 4, 16]. More generally, this technique of uncrossing to obtain a structured family of tight constraints has been utilized for various combinatorial-optimization problems; see, e.g., [18].
As mentioned earlier, the main complication with $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ is that the underlying cut-requirement function $g^{\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}}$ is usually not weakly-supermodular, even if the base cut-requirement function $f$ is. In particular, the cut-requirement function $g^{\mathsf{CR}\text{-}\mathsf{SNDP}}$ defining cut-relative $\mathsf{SNDP}$ is not weakly-supermodular. This difficulty was also noted by Dinitz et al. [8] when they considered $k$-edge fault-tolerant subgraph ($k\text{-}\mathsf{EFTS}$, or path- or cut- relative $k\text{-}\mathsf{ECSS}$ in our terminology). For this special case, they identify a notion called local weak-supermodularity that is satisfied by the cut-requirement function $\min\{k,|\delta_{G}(S)\}$ for $k\text{-}\mathsf{EFTS}$, and observed that Jain’s machinery can be used as is if this property holds; that is, one can uncross tight sets to obtain a laminar family, and hence obtain a $2$-approximation for $k\text{-}\mathsf{EFTS}$ proceeding as in Jain’s algorithm for $\mathsf{SNDP}$.
However, for general $\mathsf{CR}\text{-}\mathsf{SNDP}$, such an approach breaks down, because, as we show in Appendix 2.1, one can construct quite simple $\mathsf{CR}\text{-}\mathsf{SNDP}$ instances where an (optimal) extreme point cannot be defined by any laminar family of tight cut constraints. In particular, this implies that the cut-requirement function $g^{\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}}$ is (in general) not locally weakly-supermodular (even when $f$ is weakly-supermodular); moreover, it does not satisfy any property that allows for the uncrossing of tight sets. The upshot here is that the standard technique of uncrossing tight sets to obtain a laminar family does not work, and one needs new ideas to deal with cut-relative network design.
The chief technical novelty of our work is to show how to overcome the impediment posed by the lack of any such nice structural property for $g^{\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}}$, via a novel decomposition technique that allows one to reduce cut-relative ${f}\text{-}\mathsf{NDP}$ with a weakly supermodular base cut-requirement function to (standard) ${f}\text{-}\mathsf{NDP}$ with (different) weakly supermodular cut-requirement functions (see Theorem 4.1).
We introduce a little notation to describe this. Let $\mathsf{CR}\text{-}{(h,D)}\text{-}\mathsf{NDP}$ denote $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ on graph $D$ with base cut-requirement function $h$. For $S\subseteq V$, define its deficiency to be $f(S)-|\delta_{G}(S)|$; we say that $S$ is a small cut if its deficiency is positive (i.e., $|\delta_{G}(S)|<f(S)$). We use $G[{S}]$ to denote the subgraph induced by $S$. Let $\overline{S}$ denote $V-S$.
Assume that the underlying base cut-requirement function $f$ is weakly supermodular (as is the case with $\mathsf{SNDP}$). Observe first that if there is no small cut then $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ is the same as ${f}\text{-}\mathsf{NDP}$. Our key technical insight is that, otherwise, we can utilize a novel splitting operation to simplify our problem. We show (see Theorem 4.7) that if we pick a small cut $A\subseteq V$ with maximum deficiency, then one can define suitable weakly-supermodular functions $f_{A}:2^{A}\mapsto\mathbb{Z}$ and $f_{\overline{A}}:2^{\overline{A}}\mapsto\mathbb{Z}$ such that any solution to the original $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$-instance yields solutions to $\mathsf{CR}\text{-}{(f_{A},G[{A}])}\text{-}\mathsf{NDP}$ and $\mathsf{CR}\text{-}{(f_{\overline{A}},G[{\overline{A}}])}\text{-}\mathsf{NDP}$, and vice-versa. The intuition here is as follows. Since $A$ is a small cut, any feasible $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ solution must include all edges in $\delta_{G}(A)$. The functions $f_{A}$ and $f_{\overline{A}}$ are essentially the restrictions of $f$ to $A$ and $\overline{A}$ respectively, taking this into account. Consequently, moving to $\mathsf{CR}\text{-}{(f_{A},G[{A}])}\text{-}\mathsf{NDP}$ and $\mathsf{CR}\text{-}{(f_{\overline{A}},G[{\overline{A}}])}\text{-}\mathsf{NDP}$ does not really impact the constraints for cuts that do not cross $A$; but we do eliminate the constraints for crossing cuts. This is the simplification that we obtain via the splitting operation, and one needs to argue that this does not hurt feasibility; that is, the constraints for such cuts are still satisfied when we take $\delta_{G}(A)$ along with the $\mathsf{CR}\text{-}{(f_{A},G[{A}])}\text{-}\mathsf{NDP}$ and $\mathsf{CR}\text{-}{(f_{\overline{A}},G[{\overline{A}}])}\text{-}\mathsf{NDP}$ solutions. Here we exploit the weak-supermodularity of $f$, which implies that the deficiency function $\operatorname{def}(S):=f(S)-|\delta_{G}(S)|$ is also weakly supermodular. One can use this, and the fact that $A$ has maximum deficiency, to bound the deficiency of any crossing set $T$ in terms of the deficiency of non-crossing sets, and hence show that the cut-constraint for $T$ is satisfied.
By recursively applying the above splitting operation to $A$ and $\overline{A}$, we end up with a partition $V_{1},\ldots,V_{k}$ of $V$ and weakly-supermodular functions $f_{i}:2^{V_{i}}\mapsto\mathbb{Z}$ for all $i\in[k]$ such that, for each $i\in[k]$, the graph $G[{V_{i}}]$ contains no small cut with respect to $f_{i}$. It follows that solving ${f_{i}}\text{-}\mathsf{NDP}$ on the graph $G[{V_{i}}]$ for all $i\in[k]$ yields a solution to the original $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ instance (Theorem 4.1).
While we can obtain this decomposition efficiently, assuming that one has suitable algorithmic primitives involving the base cut-requirement function $f$ (see Section 6), in fact we only need this decomposition in the analysis: in order to obtain the stated LP-relative $2$-approximation algorithm, we only need to be able to find an extreme-point optimal solution to the natural LP-relaxation LABEL:crlp for $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ (also after fixing some variables to $1$). This is because the translation between $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$-solutions for $G$, and ${f_{i}}\text{-}\mathsf{NDP}$-solutions on $G[{V_{i}}]$ for all $i\in[k]$, also applies to fractional solutions, and implies that any extreme-point solution to LABEL:crlp yields extreme-point solutions to the LP-relaxations of ${f_{i}}\text{-}\mathsf{NDP}$ on $G[{V_{i}}]$ for all $i\in[k]$ (parts (b) and (c) of Theorem 4.1); hence, by Jain’s result, we can always find some fractional edge of value at least $\frac{1}{2}$, round its value to $1$, and iterate. For $\mathsf{SNDP}$, we can find an extreme-point optimal solution efficiently because we argue that one can devise an efficient separation oracle for LABEL:crlp (Lemma 2.3). This yields a $2$-approximation algorithm for $\mathsf{CR}\text{-}\mathsf{SNDP}$, and more generally $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$, whenever $f$ is weakly supermodular and LABEL:crlp can be solved efficiently (see Algorithm CRNDP-Alg in Section 5).
Other related work.
Network design is a fairly broad research topic, with a vast amount of literature. We limit ourselves to the work that is most closely connected to our work.
As mentioned earlier, path-relative $\mathsf{SNDP}$ was introduced by Dinitz et al. [8]. In addition to $k\text{-}\mathsf{EFTS}$, which is the path-relative version of $k\text{-}\mathsf{ECSS}$, they study the path-relative versions of two other special cases of $\mathsf{SNDP}$, for which they developed constant-factor approximation algorithms: (1) $\mathsf{SNDP}$ with $r_{i}\leq 2$ for all $i$; (2) one $s_{i}$-$t_{i}$ pair with $r_{i}\leq 3$. Both results were subsequently extended by [9] who gave a $2$-approximation for $\mathsf{SNDP}$ with $r_{i}\leq 3$ for all $i$, and a $2^{O(k^{2})}$-approximation when there is one $s_{i}$-$t_{i}$ pair with $r_{i}=k$. The only result known for general path-relative $\mathsf{SNDP}$ is due to [7], who (among other results) devise an $O(k^{2}\log^{2}n\log\log n)$-approximation algorithm in $n^{O(k)}$ time. As noted earlier, on the hardness side, prior to our work, it was not known if the setting with one $s_{i}$-$t_{i}$ pair is even NP-hard.
An influential paper by Goemans and Williamson [10] (see also the survey [11]), which built upon the work of [1] for the Steiner forest problem, developed the primal-dual method for network-design problems. Their work also popularized the use of cut-requirement functions to specify network-design problems, by showing how this framework can be used to capture a slew of problems, and how their primal-dual framework leads to a $2$-approximation algorithm for $\{0,1\}$-cut-requirement functions satisfying some properties. This work was extended to handle general integer-valued cut-requirement functions, which captures $\mathsf{SNDP}$, in [20, 12], which led to an $O(\log\max_{i}r_{i})$-approximation for $\mathsf{SNDP}$. The seminal work of Jain [15] later improved this to a $2$-approximation, and in doing so, introduced the powerful technique of iterative rounding. Later work of [17, 19] on degree-bounded network design added another ingredient to iterative rounding, namely iteratively dropping some constraints. This paradigm of iterative rounding and relaxation has proved to be an extremely versatile and powerful tool in developing algorithms (and structural results) for network-design problems, and more generally in combinatorial optimization; see [18] for an extensive study.
Iterative rounding and relaxation derives its power from the fact that an extreme-point solution to an LP-relaxation of the problem can be defined using a structured family of tight constraints, such as, often a laminar family of cut constraints in the case of network-design problems. As noted earlier, this key property does not hold for cut-relative $\mathsf{SNDP}$ (and path-relative $\mathsf{SNDP}$). Recently, various works have considered network-design problems that give rise to cut constraints that are less structured and not quite uncrossable; see [Bansal25, 3, 6] and the references therein. Our work can be seen as furthering this research direction.
2 Preliminaries and notation
Recall that we are given an undirected graph $G=(V,E)$ with nonnegative edge costs $\{c_{e}\}_{e\in E}$. For a subset $S\subseteq V$ and subset $Z\subseteq E$, which we will interchangeably view as the subgraph $(V,Z)$ of $G$, we use $\delta_{Z}(S)$ to denote $\delta_{G}(S)\cap Z$, i.e., the edges of $Z$ on the boundary of $S$. Recall that $\overline{S}$ denotes $V-S$, and $G[{S}]$ denotes the subgraph induced by $S$. A cut-requirement function on $G$ is a function $f:2^{V}\mapsto\mathbb{Z}$. (We allow $f(S)$ to be negative chiefly for notational simplicity; this does not affect anything, as we will only examine sets for which $f(S)\geq 0$.) Recall the following network-design problems.
  1. $\bullet$
    $f$-network-design problem (${f}\text{-}\mathsf{NDP}$): find a min-cost edge-set $H$ such that $|\delta_{H}(S)|\geq f(S)$ for all $S\subseteq V$.
  2. $\bullet$
    Cut-relative network-design problem ($\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$): find a min-cost edge-set $H$ such that $|\delta_{H}(S)|\geq g^{\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}}(S):=\min\bigl{\{}f(S),|\delta_{G}(S)|\bigr{\}}$ for all $S\subseteq V$. This corresponds to ${g^{\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}}}\text{-}\mathsf{NDP}$. We call $f$ the base cut-requirement function.
  3. $\bullet$
    Survivable network design problem ($\mathsf{SNDP}$) is a special case of ${f}\text{-}\mathsf{NDP}$. The input also specifies $k$ tuples $\{(s_{i},t_{i},r_{i})\}_{i\in[k]}$, where $s_{i},t_{i}\in V$ and $r_{i}\in\mathbb{Z}_{+}$ for all $i\in[k]$. Defining $f^{\mathsf{SNDP}}(S):=\max\,\{r_{i}:|S\cap\{s_{i},t_{i}\}|=1\}$ for all $S\subseteq V$, we obtain that $\mathsf{SNDP}$ is the $f^{\mathsf{SNDP}}$-network-design problem.
  4. $\bullet$
    Cut-relative $\mathsf{SNDP}$ ($\mathsf{CR}\text{-}\mathsf{SNDP}$): The input here is the same as in $\mathsf{SNDP}$. In Definition 1.1 (R2), $\mathsf{CR}\text{-}\mathsf{SNDP}$ was defined as: find a min-cost $H\subseteq E$ such that for every $i\in[k]$, every $F\subseteq E$ with $|F|<r_{i}$, and every $s_{i}$-$t_{i}$ cut $S$, we have $\delta_{H-F}(S)\neq\emptyset$ whenever $\delta_{G-F}(S)\neq\emptyset$. Later in Section 1, we stated that this can be equivalently formulated as $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ with base cut-requirement function $f^{\mathsf{SNDP}}$. We now justify this statement.
    Suppose that $H$ is feasible under (R2) but $|\delta_{H}(S)|<\min\bigl{\{}f^{\mathsf{SNDP}}(S),|\delta_{G}(S)|\bigr{\}}$ for some $S\subseteq V$. Suppose $f^{\mathsf{SNDP}}(S)=r_{i}$. Then taking $F=\delta_{H}(S)$, we have $|F|<r_{i}$, $\delta_{G-F}(S)\neq\emptyset$, but $\delta_{H-F}(S)=\emptyset$, yielding a contradiction. So $H$ is feasible for $\mathsf{CR}\text{-}{f^{\mathsf{SNDP}}}\text{-}\mathsf{NDP}$. Conversely, suppose $H$ is feasible for $\mathsf{CR}\text{-}{f^{\mathsf{SNDP}}}\text{-}\mathsf{NDP}$. Consider any $i\in[k]$, any $F\subseteq E$ with $|F|<r_{i}$, and any $s_{i}$-$t_{i}$ cut $S$ with $\delta_{G-F}(S)\neq\emptyset$. Then, $|\delta_{F}(S)|<\min\bigl{\{}r_{i},|\delta_{G}(S)|\bigr{\}}\leq\min\bigl{\{}f(S),|\delta_{G}(S)|\bigr{\}}\leq|\delta_{H}(S)|$; so $\delta_{H-F}(S)\neq\emptyset$, showing that $H$ satisfies (R2).
We will consider network-design problems defined on various graphs and with various cut-requirement functions. Given a graph $D=(V_{D},E_{D})$ and cut-requirement function $h:2^{V_{D}}\mapsto\mathbb{Z}$, we use ${(h,D)}\text{-}\mathsf{NDP}$ to denote ${f}\text{-}\mathsf{NDP}$ on the graph $D$ with cut-requirement function $h$, and $\mathsf{CR}\text{-}{(h,D)}\text{-}\mathsf{NDP}$ to denote $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ on the graph $D$ with base cut-requirement function $h$.
Properties of cut-requirement functions.
Let $f:2^{V}\mapsto\mathbb{R}$. We say that $f$ is weakly supermodular if for any two node-sets $A,B\subseteq V$, we have $f(A)+f(B)\leq\max\bigl{\{}f(A\cap B)+f(A\cup B),\,f(A-B)+f(B-A)\bigr{\}}$. We say that $f$ is symmetric if $f(S)=f(V-S)$ for all $S\subseteq V$, and is normalized if $f(\emptyset)=f(V)=0$.
Since we are working on undirected graphs, we can always replace any cut-requirement function $f$ by its symmetric, normalized version $f^{\mathsf{sym}}$, where $f^{\mathsf{sym}}(S):=\max\{f(S),f(V-S)\}$ for all $\emptyset\neq S\subsetneq V$ and $f^{\mathsf{sym}}(\emptyset)=f^{\mathsf{sym}}(V):=0$, without changing the resulting network-design problem(s), i.e., we have ${f}\text{-}\mathsf{NDP}\equiv{f^{\mathsf{sym}}}\text{-}\mathsf{NDP}$ and $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}\equiv\mathsf{CR}\text{-}{f^{\mathsf{sym}}}\text{-}\mathsf{NDP}$.
Lemma 2.1 shows that weak-supermodularity is preserved under various operations. This will be quite useful in the analysis of our decomposition technique. Part (a) of the lemma shows that weak supermodularity is preserved under symmetrization and normalization, so for any ${f}\text{-}\mathsf{NDP}$ or $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ instance involving a weakly-supermodular base cut-requirement function $f$, we may assume that $f$ is symmetric and normalized.
Lemma 2.1.
Let $f:2^{V}\mapsto\mathbb{R}$ be weakly supermodular. Then
  1. (a)
    $f^{\mathsf{sym}}$ is weakly supermodular.
  2. (b)
    Let $w\in\mathbb{R}_{+}^{E}$. Define $g_{1}(S):=f(S)-w\bigl{(}\delta_{G}(S)\bigr{)}$ for all $S\subseteq V$. Then $g_{1}$ is weakly supermodular, and $g_{1}$ is symmetric if $f$ is symmetric. In particular, for any $Z\subseteq E$, the function $f(S)-|\delta_{Z}(S)|$ is weakly supermodular, and is symmetric if $f$ is symmetric.
  3. (c)
    Let $X\subseteq V$, and $\mathcal{C}$ be a collection of subsets of $V-X$ closed under taking set intersections, unions, differences, and complements with respect to $V-X$, i.e., if $S\in\mathcal{C}$ then $V-X-S\in\mathcal{C}$, and $S,T\in\mathcal{C}$ implies that $\{S\cap T,S\cup T,S-T,T-S\}\subseteq\mathcal{C}$.
    Define the projection of $f$ onto $X$ with respect to $\mathcal{C}$ as the function $g_{2}:2^{X}\mapsto\mathbb{R}$ given by $g_{2}(\emptyset)=g_{2}(X):=0$, and $g_{2}(S):=\max_{T\in\mathcal{C}}f(S\cup T)$ for all $\emptyset\neq S\subsetneq X$. If $f$ is symmetric, then $g_{2}$ is weakly supermodular and symmetric.
Proof 2.2.
Part (a) follows from a simple case analysis; part (b) follows because $w\bigl{(}\delta_{G}(S)\bigr{)}$ is symmetric, submodular. For part (c), the closure properties of $\mathcal{C}$ enable us to transfer the weak-supermodularity of $f$ to the function $g_{2}$.
For part (a), consider any $A,B\subseteq V$. If $A\subseteq B$ or $B\subseteq A$, then $f^{\mathsf{sym}}(A)+f^{\mathsf{sym}}(B)=f^{\mathsf{sym}}(A\cap B)+f^{\mathsf{sym}}(A\cup B)$; if $A\cap B=\emptyset$ or $V-(A\cup B)=\emptyset$, then $\{A-B,B-A\}=\{A,B\}$ or $\{A-B,B-A\}=\{V-B,V-A\}$, so $f^{\mathsf{sym}}(A)+f^{\mathsf{sym}}(B)=f^{\mathsf{sym}}(A-B)+f^{\mathsf{sym}}(B-A)$, where we also utilize the symmetry of $f^{\mathsf{sym}}$.
So assume that $A-B,B-A,A\cap B,V-(A\cup B)$ are all non-empty. We have $f^{\mathsf{sym}}(A)=f(X)$, where $X\in\{A,V-A\}$, and $f^{\mathsf{sym}}(B)=f(Y)$, where $Y\in\{B,V-B\}$. So
$\begin{split}f^{\mathsf{sym}}(A)+f^{\mathsf{sym}}(B)&\leq\max\bigl{\{}f(X\cap Y)+f(X\cup Y),\,f(X-Y)+f(Y-X)\bigr{\}}\\ &\leq\max\bigl{\{}f^{\mathsf{sym}}(X\cap Y)+f^{\mathsf{sym}}(X\cup Y),\ f^{\mathsf{sym}}(X-Y)+f^{\mathsf{sym}}(Y-X)\bigr{\}}.\end{split}$
We show that
$\begin{split}f^{\mathsf{sym}}(X\cap Y)+f^{\mathsf{sym}}(X\cup Y)&\in\bigl{\{}f^{\mathsf{sym}}(A\cap B)+f^{\mathsf{sym}}(A\cup B),\ f^{\mathsf{sym}}(A-B)+f^{\mathsf{sym}}(B-A)\bigr{\}},\\ f^{\mathsf{sym}}(X-Y)+f^{\mathsf{sym}}(Y-X)&\in\bigl{\{}f^{\mathsf{sym}}(A\cap B)+f^{\mathsf{sym}}(A\cup B),\ f^{\mathsf{sym}}(A-B)+f^{\mathsf{sym}}(B-A)\bigr{\}}.\end{split}$ (1)
which will complete the proof of part (a). If $X=A$ and $Y=B$, then (1) clearly holds. So consider the remaining cases.
  1. $\bullet$
    $X=A$,  $Y=V-B$. We have $X\cap Y=A-B$, and $X\cup Y=V-(B-A)$, so $f^{\mathsf{sym}}(X\cup Y)=f^{\mathsf{sym}}(B-A)$. Also, $X-Y=A\cap B$, and $Y-X=V-(A\cup B)$, so $f^{\mathsf{sym}}(Y-X)=f^{\mathsf{sym}}(A\cup B)$. So (1) holds.
  2. $\bullet$
    $X=V-A$,  $Y=B$. This is completely symmetric to the above case. We have $f^{\mathsf{sym}}(X\cap Y)=f^{\mathsf{sym}}(B-A)$ and $f^{\mathsf{sym}}(X\cup Y)=f^{\mathsf{sym}}(A-B)$. Also, $f^{\mathsf{sym}}(X-Y)=f^{\mathsf{sym}}(A\cup B)$ and $f^{\mathsf{sym}}(Y-X)=f^{\mathsf{sym}}(A\cap B)$.
  3. $\bullet$
    $X=V-A$,  $Y=V-B$. We have $X\cap Y=V-(A\cup B)$, so $f^{\mathsf{sym}}(X\cap Y)=f^{\mathsf{sym}}(A\cup B)$, and $X\cup Y=V-(A\cap B)$, so $f^{\mathsf{sym}}(X\cup Y)=f^{\mathsf{sym}}(A\cap B)$. Also, $X-Y=B-A$ and $Y-X=A-B$.
For part (b), since $w\bigl{(}\delta_{G}(S)\bigr{)}$ is a symmetric, submodular function, we have
$w\bigl{(}\delta_{G}(A)\bigr{)}+w\bigl{(}\delta_{G}(B)\bigr{)}\geq\max\bigl{\{}w\bigl{(}\delta_{G}(A\cap B)\bigr{)}+w\bigl{(}\delta_{G}(A\cup B)\bigr{)},\ w\bigl{(}\delta_{G}(A-B)\bigr{)}+w\bigl{(}\delta_{G}(B-A)\bigr{)}\bigr{\}}$
for all $A,B\subseteq V$. Coupled with the weak-supermodularity of $f$, this yields that $g_{1}$ is weakly supermodular, and is symmetric if $f$ is symmetric.
Finally, consider part (c). We first argue that $g_{2}$ is symmetric. If $A\in\{\emptyset,X\}$, we have $g_{2}(A)=g_{2}(X-A)$ by definition. Otherwise, if $g_{2}(A)=f(A\cup T_{A})$, we have $f(A\cup T_{A})=f\bigl{(}V-(A\cup T_{A})\bigr{)}=f\bigl{(}(X-A)\cup(V-X-T_{A})\bigr{)}\leq g_{2}(X-A)$ where the last inequality follows since $V-X-T_{A}\in\mathcal{C}$. We also have $g_{2}(X-A)\leq g_{2}(A)$ by the same type of reasoning, and so $g_{2}(A)=g_{2}(X-A)$ for all $A\subseteq X$.
Now we show that $g_{2}$ is weakly supermodular. Consider $A,B\subseteq X$. If any of the sets $A-B,B-A,A\cap B,X-(A\cup B)$ are empty, then $g_{2}(A)+g_{2}(B)=g_{2}(A\cap B)+g_{2}(A\cup B)$ or $g_{2}(A)+g_{2}(B)=g_{2}(A-B)+g_{2}(B-A)$. The former holds when $A-B$ or $B-A$ is empty; the latter holds when $A\cap B$ or $X-(A\cup B)$ is empty, where we also utilize the fact that $g_{2}$ is symmetric. So assume otherwise. Then, for every $S\in\{A,B,A\cap B,A\cup B,A-B,B-A\}$, we have $g_{2}(S)=\max_{T\in\mathcal{C}}f(S\cup T)$.
Let $g_{2}(A)=f(A\cup T_{A})$ and $g_{2}(B)=f(B\cup T_{B})$, where $T_{A},T_{B}\in\mathcal{C}$. Let $A^{\prime}=A\cup T_{A}$ and $B^{\prime}=B\cup T_{B}$. For any $S\in\{A^{\prime}\cap B^{\prime},A^{\prime}\cup B^{\prime},A^{\prime}-B^{\prime},B^{\prime}-A^{\prime}\}$, we can write $S=(S\cap X)\cup T$, where $T\in\mathcal{C}$ due to the closure properties of $\mathcal{C}$. It follows that $g_{2}(S\cap X)\geq f(S)$ for all $S\in\{A^{\prime}\cap B^{\prime},A^{\prime}\cup B^{\prime},A^{\prime}-B^{\prime},B^{\prime}-A^{\prime}\}$. So if $f(A^{\prime})+f(B^{\prime})\leq f(A^{\prime}\cap B^{\prime})+f(A^{\prime}\cup B^{\prime})$, we obtain that $g_{2}(A)+g_{2}(B)\leq g_{2}(A\cap B)+g_{2}(A\cup B)$, and if $f(A^{\prime})+f(B^{\prime})\leq f(A^{\prime}-B^{\prime})+f(B^{\prime}-A^{\prime})$, we obtain that $g_{2}(A)+g_{@}(B)\leq g_{2}(A-B)+g_{2}(B-A)$.
The intuition behind the projection operation is as follows. Suppose $f$ is a cut-requirement function and we know that no edges are picked from $E_{G}(V-X)\cup\delta_{G}(X)$. Then, roughly speaking, $g_{2}(S)$ captures the constraints that arise on $S\subseteq X$ due to $f$ and $\mathcal{C}$.
2.1 LP-relaxation for $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$
We consider the following natural LP-relaxation for $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$. From now on, we assume that the cut-requirement function $f:2^{V}\mapsto\mathbb{Z}$ is weakly supermodular, symmetric, and normalized. Given $\alpha\in\mathbb{R}^{E}$ and $F\subseteq E$, let $\alpha(F)$ denote $\sum_{e\in F}\alpha_{e}$.
$\min$ $\sum_{e\in E}c_{e}x_{e}$ $(\text{{CRLP}}_{f,G})$
s.t. $x\bigl{(}\delta_{G}(S)\bigr{)}$ $\geq\min\bigl{\{}f(S),|\delta_{G}(S)|\bigr{\}}\qquad$ $\forall S\subseteq V$ (2)
$0\leq x_{e}$ $\leq 1$ $\forall e\in E.$ (3)
The LP-relaxation for ${f}\text{-}\mathsf{NDP}$ is similar, with the cut-constraints taking the simpler form $x\bigl{(}\delta_{G}(S)\bigr{)}\geq f(S)$ for all $S\subseteq V$. We denote this LP by $(\text{{NDP}}_{f,G})$. Our algorithm is based on iterative rounding, where we iteratively fix some $x_{e}$ variables to $1$. Suppose that we have fixed $x_{e}=1$ for all $e\in Z_{1}$. Then, in that iteration, we are considering the residual problem on the subgraph $G^{\prime}=(V,E^{\prime}=E-Z_{1})$, which leads to the following variant of $(\text{{CRLP}}_{f,G})$:
$\min$ $\sum_{e\in E^{\prime}}c_{e}x_{e}$ (P’)
s.t. $x\bigl{(}\delta_{G^{\prime}}(S)\bigr{)}$ $\geq\min\bigl{\{}f(S),|\delta_{G}(S)|\bigr{\}}-|\delta_{Z_{1}}(S)|\qquad$ $\forall S\subseteq V$ (4)
$0\leq x_{e}$ $\leq 1$ $\forall e\in E^{\prime}.$
Defining $f^{\prime}(S)=f(S)-|\delta_{Z_{1}}(S)|$ for all $S\subseteq V$, we can rephrase constraints (4) as $x\bigl{(}\delta_{G^{\prime}}(S)\bigr{)}\geq\min\bigl{\{}f^{\prime}(S),|\delta_{G^{\prime}}(S)|\}$ for all $S\subseteq V$. Since $f^{\prime}$ is weakly supermodular, symmetric (Lemma 2.1 (b)), and normalized, (P’) is of the same form as $(\text{{CRLP}}_{f,G})$: we have (P’) $\equiv(\text{{CRLP}}_{f^{\prime},G^{\prime}})$. This will be convenient for iterative rounding, as the residual problem is of the same form as the original problem. $(\text{{CRLP}}_{f,G})$ and (P’) involve an exponential number of constraints, but using the ellipsoid method, we can find an extreme-point optimal solution provided we have a separation oracle for them. Recall that a separation oracle is a procedure that given a candidate point determines if it is feasible, or returns a violated inequality. Observe that a separation oracle for $(\text{{CRLP}}_{f,G})$ also yields a separation oracle for (P’), since (P’) is simply $(\text{{CRLP}}_{f,G})$ after fixing some variables to $1$.
For $\mathsf{CR}\text{-}\mathsf{SNDP}$, i.e., when $f=f^{\mathsf{SNDP}}$, a polytime separation oracle for $(\text{{CRLP}}_{f,G})$ can be obtained using min-cut computations. This was shown by [8] for path-relative $k\text{-}\mathsf{ECSS}$, which is equivalent to cut-relative $k\text{-}\mathsf{ECSS}$ (Lemma A.1); refining their ideas yields a separation oracle for $(\text{{CRLP}}_{f,G})$ for general $\mathsf{SNDP}$.
Lemma 2.3.
When $f=f^{\mathsf{SNDP}}$, there is a polytime separation oracle for $(\text{{CRLP}}_{f,G})$. Hence, one can obtain extreme-point optimal solutions to $(\text{{CRLP}}_{f,G})$ and (P’) in polytime.
Proof 2.4.
The second statement follows from Fact 7. Let $(s_{i},t_{i},r_{i})_{i\in[k]}$ be the $k$ tuples defining the $\mathsf{SNDP}$ instance, which recall leads to the cut-requirement function $f^{\mathsf{SNDP}}(S):=\max\,\{r_{i}:|S\cap\{s_{i},t_{i}\}|=1\}$ for all $S\subseteq V$. Say that $S\subseteq V$ is a small cut if $f^{\mathsf{SNDP}}(S)>|\delta_{G}(S)|$, and it is a small cut for $i$ if $|S\cap\{s_{i},t_{i}\}|=1$ and $r_{i}>|\delta_{G}(S)|$. Note that if $e\in\delta_{G}(S)$ for any small cut $S$, then we must have $x_{e}=1$ in any feasible solution.
Now we can equivalently phrase the constraints of $(\text{{CRLP}}_{f,G})$ as follows: for any edge $a=uv\in E$, we must have $x_{a}=1$ or, for every $i\in[k]$, the minimum $s_{i}$-$t_{i}$ cut also separating $u$ and $v$ under the $\{x_{e}\}_{e\in E}$-capacities should have capacity at least $r_{i}$. So to detect if $x$ is a feasible solution, we simply need to consider every edge $a=uv$ with $x_{a}<1$ and every $i\in[k]$, and find the minimum $s_{i}$-$t_{i}$ cut also separating $u$ and $v$ under the $\{x_{e}\}_{e\in E}$ capacities. If an $s_{i}$-$t_{i}$ cut separates $u,v$, then $u$ lies on the $s_{i}$-side and $v$ lies on the $t_{i}$-side, or the other way around. So to find the minimum $s_{i}$-$t_{i}$ cut also separating $u$ and $v$, we find the minimum $\{s_{i},u\}$-$\{t_{i},v\}$ cut—by which we mean the minimum $s_{i}$-$t_{i}$ cut where $s_{i},u$ are on the same side, and $t_{i},v$ are on the other side–and the minimum $\{s_{i},v\}$-$\{t_{i},u\}$-cut and take the one with smaller capacity.
Extreme-point solution to $(\text{{CRLP}}_{f,G})$ not defined by a laminar family
We show that for the $\mathsf{CR}\text{-}\mathsf{SNDP}$ instance shown in Fig. 1, the extreme-point solution $\widehat{x}$ shown in the figure, cannot be expressed as the solution to any laminar family of tight cut constraints (2) (along with tight edge constraints (3)). The $\mathsf{SNDP}$ instance is $s$-$t$ $2$-edge connectivity: that is, the base cut-requirement function is given by $f(S)=2$ if $S$ is an $s$-$t$ cut, and $f(S)=0$ otherwise.
Figure 1: $\mathsf{CR}\text{-}\mathsf{SNDP}$ instance, and an extreme-point solution $\widehat{x}$ to $(\text{{CRLP}}_{f,G})$. The solid edges have $\widehat{x}_{e}=1$ and the dashed edges have $\widehat{x}_{e}=0.5$, so $\widehat{x}_{e}=0.5$ for $e\in\{su,uw,sv,vw\}$ and $\widehat{x}_{e}=1$ for $e\in\{sw,wt\}$.
Note that the only small cut above containing $s$ is $S=\{s,u,v,w\}$. We argue that: (1) $\widehat{x}$ is an extreme-point solution to $(\text{{CRLP}}_{f,G})$; and (2) $\widehat{x}$ cannot be defined using a laminar family of cut constraints.
$\widehat{x}$ is an (optimal) extreme point.
It suffices to show that $\widehat{x}$ is feasible, and that it is the unique solution to a collection of tight constraints. It is straightforward to verify feasibility of $\widehat{x}$ by inspection. Consider the following set of equations, which come from constraints that are tight at $\widehat{x}$.
$x\bigl{(}\delta(S)\bigr{)}$ $=2\qquad\forall S\in\bigl{\{}\{s\},\{s,u,v\},\{s,u,w\},\{s,v,w\}\bigr{\}},$ (5)
$x_{e}$ $=1\qquad\forall e\in\{sw,wt\}.$ (6)
We claim that $\widehat{x}$ is the unique solution to these equations, showing that $\widehat{x}$ is an extreme-point solution to $(\text{{CRLP}}_{f,G})$. Indeed, after substituting $x_{sw}=x_{wt}=1$, by considering every pair of sets in (5), we obtain that all edges in $\{su,uw,sv,vw\}$ must have the same $x_{e}$ value. (For example, $x\bigl{(}\delta(s)\bigr{)}=2=x\bigl{(}\delta(\{s,u,w\})\bigr{)}$ implies that $x_{vw}=x_{su}$.) Together with $x\bigl{(}\delta(s)\bigr{)}=2$ and $x_{sw}=1$, this fixes the values of all edges in $\{su,uw,sv,vw\}$ to $0.5$.
Taking edge costs $c_{sw}=0$ and $c_{e}=1$ for all other edges, it is not hard to see that $\widehat{x}$ is an optimal solution. The constraints $x\bigl{(}\delta(S)\bigr{)}\geq 2$ for all $S\in\bigl{\{}\{s\},\{s,u,v\},\{s,u,w\},\{s,v,w\}\bigr{\}}$ imply that (since $x_{e}\leq 1$ for all $e$)
$x_{su}+x_{sv}\geq 1,\quad x\bigl{(}\delta(u)\bigr{)}\geq 1,\quad x\bigl{(}\delta(v)\bigr{)}\geq 1,\quad x_{uw}+x_{vw}\geq 1.$
Adding these constraints, we obtain that $x_{su}+x_{sv}+x_{uv}+x_{vw}\geq 2$. Therefore, the cost of any feasible solution is at least $3$, and we have $c^{T}\widehat{x}=3$.
$\widehat{x}$ cannot be defined via a laminar family.
We now argue that $\widehat{x}$ cannot be defined by any laminar family of tight cut constraints and tight edge constraints. Note that any tight cut constraint corresponds to an $s$-$t$ cut. So a laminar family $\mathcal{L}$ of tight cut constraints consists of a union of two chains $\mathcal{L}_{1}\cup\mathcal{L}_{2}$, where a chain is a nested collection of sets: $\mathcal{L}_{1}$ comprising sets containing $s$ but not $t$, and $\mathcal{L}_{2}$ comprising sets containing $t$ but not $s$. Then $\mathcal{L}^{\prime}=\mathcal{L}_{1}\cup\{V-S:S\in\mathcal{L}_{2}\}$ is also a laminar family, yielding the same cut-constraints as $\mathcal{L}$, and and all sets in $\mathcal{L}^{\prime}$contain $s$ but not $t$. So we only need to consider laminar families of tight cut constraints corresponding to a chain of sets containing $s$ and not $t$.
A simple argument now rules out the existence of any such chain that can uniquely define $\widehat{x}$ along with the tight edge constraints (6). Any such chain of sets $\mathcal{C}$ is obtained by ordering the elements $u,v,w$, and taking sets of the form $\{s\}\cup T$, where $T$ is some prefix (including the empty prefix) of this ordered sequence; so $\mathcal{C}$ contains at most $4$ sets. Since $\widehat{x}$ has $4$ fractional edges, $E^{\prime}=\{su,sv,uw,vw\}$, it must be that $\bigl{\{}\chi^{\delta_{E^{\prime}}(S)}\}_{S\in\mathcal{C}}$ contains $4$ linearly independent vectors. So $\mathcal{C}$ contains exactly $4$ sets, and in particular the set $\{s,u,v,w\}$, and all vectors in $\bigl{\{}\chi^{\delta_{E^{\prime}}(S)}\}_{S\in\mathcal{C}}$ are linearly independent. This yields a contradiction, since $\chi^{\delta_{E^{\prime}}(\{s,u,v,w\})}=\vec{0}$.
3 Modeling power of $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$
We prove that the $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ examples listed under “Modeling power” in Section 1 (Examples 1.2 and 1.3) model the stated problems. Recall that in both examples, we have a non-increasing function $\tau:\mathbb{Z}_{+}\mapsto\mathbb{Z}_{+}$.
Theorem 3.1.
Consider the base cut-requirement function $f^{\mathsf{grace}}$ in Example 1.2, where $f^{\mathsf{grace}}(S):=\min\,\{\ell:\tau(\ell)<|S|\}$, for $S\neq\emptyset$.
  1. (a)
    The function $f^{\mathsf{grace}}$ is weakly-supermodular.
  2. (b)
    $H$ is feasible for $\mathsf{CR}\text{-}{f^{\mathsf{grace}}}\text{-}\mathsf{NDP}$ iff for every fault-set $F\subseteq E$, and every $S\subseteq V$ with $|S|\leq\tau(|F|)$, $S$ is (the node-set of) a component of $G-F$ iff it is a component of $H-F$.
Proof 3.2.
The function $f^{\mathsf{grace}}$ is downwards-monotone: if $\emptyset\neq T\subseteq S$, then $f^{\mathsf{grace}}(T)\geq f^{\mathsf{grace}}(S)$. This follows from the definition, since $\tau$ is a non-increasing function. Weak-supermodularity of $f^{\mathsf{grace}}$ is now immediate since we have $f^{\mathsf{grace}}(A)+f^{\mathsf{grace}}(B)\leq f^{\mathsf{grace}}(A-B)+f^{\mathsf{grace}}(B-A)$, if $A-B,B-A\neq\emptyset$, and otherwise $f^{\mathsf{grace}}(A)+f^{\mathsf{grace}}(B)=f^{\mathsf{grace}}(A\cap B)+f^{\mathsf{grace}}(A\cup B)$.
Note that $f^{\mathsf{grace}}$ is not symmetric or normalized, but as indicated by Lemma 2.1, we can consider $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ with the symmetric, normalized version of $f^{\mathsf{grace}}$ as the base cut-requirement function, without changing the problem.
To prove the feasibility characterization in part (b), we note that the function $f^{\mathsf{grace}}$ is constructed so that for any $S\subseteq V$ and $F\subseteq E$, we have $|S|\leq\tau(|F|)$ iff $f^{\mathsf{grace}}(S)>|F|$, so that $|\delta_{H}(S)|\geq f^{\mathsf{grace}}(S)$ implies that $S$ is not a component of $H-F$. Suppose $H$ is such that $|\delta_{H}(S)|\geq\min\bigl{\{}f^{\mathsf{grace}}(S),|\delta_{G}(S)|\bigr{\}}$ for all $S\subseteq V$. Consider any fault-set $F\subseteq E$ and any $S\subseteq V$ such that $|S|\leq\tau(|F|)$. By the above, if $S$ is a component of $H-F$, then $|\delta_{H}(S)|<f^{\mathsf{grace}}(S)$, and so $S$ must also be a component of $G-F$. But this also implies that if $S$ is a component of $G-F$, then it is a component of $H-F$; otherwise, there is some $A\subsetneq S$ that is a component of $H-F$ with $|A|<|S|\leq\tau(|F|)$, and so $A$ must be a component of $G-F$, which is not the case.
Conversely, suppose that for every $F\subseteq E$, and $S\subseteq V$ with $|S|\leq\tau(|F|)$, we have that $S$ is a component of $G-F$ iff it is a component of $H-F$. If $|\delta_{H}(S)|<\min\bigl{\{}f^{\mathsf{grace}}(S),|\delta_{G}(S)|\bigr{\}}$ for some $S\subseteq V$, then take $F=\delta_{H}(S)$. Since $f^{\mathsf{grace}}(S)>|F|$, we have $|S|\leq\tau(|F|)$. But $S$ is a component of $H-F$ and not $G-F$, which yields a contradiction.
Theorem 3.3.
Let be a monotone function, i.e., if $T\subseteq S$, as in Example 1.3. Consider $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$, where the base cut-requirement function $f$ is given by $f(S):=\min\,\{\ell:\tau(\ell)<$, for all $S\neq\emptyset$. Then
  1. (a)
    $f$ is weakly-supermodular;
  2. (b)
    $H$ is feasible for $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ iff for every fault-set $F\subseteq E$, and every $S\subseteq V$ with , $S$ is (the node-set of) a component of $G-F$ iff it is a component of $H-F$.
Proof 3.4.
We mimic the proof of Theorem 3.1. Since is monotone, we obtain that $f$ is downwards-monotone, and hence weakly-supermodular.
For part (b), similar to before, we have that iff $f(S)>|F|$. Suppose $H$ is feasible for $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$. Consider any fault-set $F\subseteq E$ and any $S\subseteq V$ be such that . By the above, $f(S)>|F|$, so if $S$ is a component of $H-F$, then it must also be a component of $G-F$. Again, this also implies that if $S$ is a component of $G-F$, then it is a component of $H-F$: otherwise, there is some $A\subsetneq S$ that is a component of $H-F$, and , which means that $A$ is a component of $G-F$.
Conversely, suppose that for every $F\subseteq E$, and $S\subseteq V$ with , we have that $S$ is a component of $G-F$ iff it is a component of $H-F$. If $|\delta_{H}(S)|<\min\bigl{\{}f(S),|\delta_{G}(S)|\bigr{\}}$ for some $S\subseteq V$, then take $F=\delta_{H}(S)$. Since $f(S)>|F|$, we have . But $S$ is a component of $H-F$ and not $G-F$, which yields a contradiction.
Our intent with Examples 1.2 and 1.3 is to illustrate that $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ can be used to model some interesting problems. The choice of the $\tau$ and/or functions can of course impact our ability to solve the LP-relaxation LABEL:crlp efficiently.
4 Structure of feasible $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ solutions: a decomposition result
We now prove our decomposition result, showing that $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ can be reduced to ${f}\text{-}\mathsf{NDP}$ with suitably-defined weakly-supermodular cut-requirement functions. Recall that the base cut-requirement function $f$ is weakly supermodular (and symmetric, normalized), but the cut-requirement function $g^{\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}}$ underlying $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ need not be. Given $x\in\mathbb{R}^{E}$ and $S\subseteq V$, we use $x^{S}$ to denote the restriction of $x$ to edges in $S$, i.e., the vector $(x_{e})_{e\in E(S)}\in\mathbb{R}^{E(S)}$.
Theorem 4.1 (Decomposition result).
There is a partition $V_{1},V_{2},\ldots,V_{k}$ of $V$ and weakly-supermodular, symmetric, and normalized functions $f_{i}:2^{V_{i}}\mapsto\mathbb{Z}$ such that the following hold.
  1. (a)
    $F\subseteq E$ is a feasible solution to $\mathsf{CR}\text{-}{(f,G)}\text{-}\mathsf{NDP}$ iff $F\supseteq\delta_{G}(V_{i})$ and $F(V_{i})$ is a feasible solution to ${(f_{i},G[{V_{i}}])}\text{-}\mathsf{NDP}$ for all $i\in[k]$.
  2. (b)
    $x\in\mathbb{R}^{E}$ is feasible to $(\text{{CRLP}}_{f,G})$ iff for all $i\in[k]$, we have $x_{e}=1$ for all $e\in\delta_{G}(V_{i})$, and $x^{V_{i}}$ is feasible to $(\text{{NDP}}_{f_{i},G[{V_{i}}]})$.
  3. (c)
    $x$ is an extreme-point solution to $(\text{{CRLP}}_{f,G})$ iff for all $i\in[k]$, we have $x_{e}=1$ for all $e\in\delta_{G}(V_{i})$, and $x^{V_{i}}$ is an extreme-point solution to $(\text{{NDP}}_{f_{i},G[{V_{i}}]})$.
Before delving into the proof of Theorem 4.1, we state the following immediate corollary of Theorem 4.1 (c), which directly leads to a $2$-approximation algorithm for $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ whenever $(\text{{CRLP}}_{f,G})$ can be solved efficiently.
Corollary 4.2.
Let $\widehat{x}$ be an extreme-point solution to $(\text{{CRLP}}_{f,G})$. There exists some $e\in E$ for which $\widehat{x}_{e}\geq 1/2$.
Proof 4.3.
By Theorem 4.1 (c), there is a partition of $V$ and ${f}\text{-}\mathsf{NDP}$-instances defined on each part such that $\widehat{x}$ induces an extreme-point solution in each ${f}\text{-}\mathsf{NDP}$-instance. By Jain’s result, this implies that in each of these instances, there is some edge $e$ with $\widehat{x}_{e}\geq 1/2$.
The rest of this section is devoted to the proof of Theorem 4.1. For a set $S\subseteq V$, define its deficiency $\operatorname{def}(S):=f(S)-|\delta_{G}(S)|$. We use $\operatorname{def}_{f,G}(S)$ (or $\operatorname{def}_{f,E}(S)$) when we want to make the cut-requirement function and graph (or edge-set) explicit. By Lemma 2.1 (b), $\operatorname{def}(S)$ is weakly supermodular; also, it is symmetric (and normalized). We say that $S$ is a small cut (or more explicitly a small-$(f,G)$-cut) if $\operatorname{def}(S)>0$. Clearly, if there is no small cut, then $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ becomes equivalent to ${f}\text{-}\mathsf{NDP}$. Our chief insight and key structural result is that if there is a small cut, then we can simplify the $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ instance and move to smaller $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ instances, by “splitting” the instance along a suitable small cut (Theorem 4.7). By repeating this operation, we eventually obtain an instance with no small cuts, which leads to Theorem 4.1. The splitting operation relies on the following modification of $f$.
Definition 4.4.
Let $\emptyset\neq S\subsetneq V$. Recall that $\overline{S}:=V-S$. Let $Z=\delta_{G}(S)$. Define the function $f_{S}:2^{S}\mapsto\mathbb{Z}$, which we call the restriction of $f$ to $S$, as follows.
$f_{S}(T)=\begin{cases}0;&\text{if $T=\emptyset$ or $T=S$}\\ \max\bigl{\{}f(T)-|\delta_{Z}(T)|,f(T\cup\overline{S})-|\delta_{Z}(T\cup\overline{S})|\bigr{\}}&\text{otherwise}\end{cases}\qquad\forall T\subseteq S.$
The intuition here is that for $T\subseteq S$, $f_{S}(T)$ is a lower bound on the residual requirement that needs to be met from $\delta_{G[{S}]}(T)$ given that we pick all edges of $Z$. Note that since $f$ and $|\delta_{Z}(.)|$ are symmetric, $f(T\cup\overline{S})-|\delta_{Z}(T\cup\overline{S})|=f(S-T)-|\delta_{Z}(S-T)|$, and so $f_{S}$ is the symmetric, normalized version of of the residual requirement function $f-|\delta_{Z}(\cdot)|$. We use the definition above, as this will be more convenient to work with later in Section 6, where we establish other useful properties of the decomposition given by Theorem 4.1.
In the splitting operation, we pick a small cut $A$ of maximum deficiency, and move to $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ instances on the smaller graphs $G[{A}]$, and $G[{\overline{A}}]$ obtained by restricting $f$ to $A$, and restricting $f$ to $\overline{A}$ respectively (Theorem 4.7). Lemma 4.5 states some basic properties of the restriction of $f$. Part (b) will be quite useful in proving that the splitting operation maintains feasibility, and part (a) allows us to repeat the splitting operation on the smaller instances. For a set $S\subseteq V$, and set $F\subseteq E$ of edges, we use $F(S)$ to denote the edges of $F$ with both ends in $S$.
Lemma 4.5.
Let $\emptyset\neq S\subsetneq V$. Let $f_{S}$ be the restriction of $f$ to $S$. Let $\delta_{G}(S)\subseteq F\subseteq E$.
  1. (a)
    $f_{S}$ is weakly supermodular, symmetric, and normalized.
  2. (b)
    Let $\emptyset\neq T\subsetneq S$. Then, $\operatorname{def}_{f_{S},F(S)}(T)=\max\bigl{\{}\operatorname{def}_{f,F}(T),\operatorname{def}_{f,F}(T\cup\overline{S})\bigr{\}}$.
Proof 4.6.
Let $Z=\delta_{G}(S)$. Define $f^{\prime}:2^{V}\mapsto\mathbb{R}$ by $f^{\prime}(T):=f(T)-|\delta_{Z}(T)|$. Then, $f^{\prime}$ is weakly supermodular and symmetric (Lemma 2.1 (b)).
For part (a), we can proceed in two ways. As noted earlier, $f_{S}$ is the symmetric, normalized version $f^{\prime}$, so part (a) follows from Lemma 2.1 (a). Alternatively, observe that $f_{S}$ is the projection of $f^{\prime}$ onto $S$ with respect to $\mathcal{C}:=\{\emptyset,\overline{S}\}$, which is closed under set intersections, unions, differences, and complements with respect to $V-S$. So by Lemma 2.1 (c), we obtain that $f_{S}$ is weakly supermodular and symmetric. Also, $f_{S}$ is normalized by definition.
For part (b), the stated equality follows by simply plugging in $f_{S}$, and noting that for $Y\in\{T,T\cup\overline{S}\}$, we have $|\delta_{F}(Y)|=|\delta_{F(S)}(T)|+|\delta_{Z}(Y)|$.
Theorem 4.7 (Splitting along a small cut).
Suppose that there is some small cut. Let $A\subseteq V$ be a maximum-deficiency cut. Note that $\operatorname{def}(A)>0$ and $\emptyset\neq A\subsetneq V$.
  1. (a)
    $F\subseteq E$ is a feasible solution to $\mathsf{CR}\text{-}{(f,G)}\text{-}\mathsf{NDP}$ iff $F\supseteq\delta_{G}(A)$, $F(A)$ is a feasible solution to $\mathsf{CR}\text{-}{(f_{A},G[{A}])}\text{-}\mathsf{NDP}$, and $F(\overline{A})$ is a feasible solution to $\mathsf{CR}\text{-}{(f_{\overline{A}},G[{\overline{A}}])}\text{-}\mathsf{NDP}$.
  2. (b)
    $x\in\mathbb{R}_{+}^{E}$ is feasible to $(\text{{CRLP}}_{f,G})$ iff $x_{e}=1$ for all $e\in\delta_{G}(A)$, $x^{A}$ is feasible to $(\text{{CRLP}}_{f_{A},G[{A}]})$, and $x^{\overline{A}}$ is feasible to $(\text{{CRLP}}_{f_{\overline{A}},G[{\overline{A}}]})$.
Proof 4.8.
Part (a) can be seen as the special case of part (b), where $x$ is integral. We prove part (a) here, as the proof is somewhat (notational) simpler. The proof of part (b) is quite similar, and is deferred to Appendix B. For notational simplicity, let $(f_{1},G_{1},F_{1})=\bigl{(}f_{A},G[{A}],F(A)\bigr{)}$ and $(f_{2},G_{2},F_{2})=\bigl{(}f_{\overline{A}},G[{\overline{A}}],F(\overline{A})\bigr{)}$. Let $Z=\delta_{G}(A)$.
The “only if” direction follows easily from the definition of the restriction of $f$. Let $F$ be feasible to $\mathsf{CR}\text{-}{(f,G)}\text{-}\mathsf{NDP}$. Since $A$ is a small cut, we must have $F\supseteq Z$. Consider any $\emptyset\neq T\subsetneq A$. Since $F$ is feasible, we have $|\delta_{F}(T)|\geq\min\bigl{\{}f(T),|\delta_{G}(T)|\bigr{\}}$, which implies that $|\delta_{F_{1}}(T)|\geq\min\bigl{\{}f(T)-|\delta_{Z}(T)|,|\delta_{G_{1}}(T)|\bigr{\}}$. Similarly, we have $|\delta_{F}(T\cup\overline{A})|\geq\min\bigl{\{}f(T\cup\overline{A}),|\delta_{G}(T\cup\overline{A})|\bigr{\}}$, so $|\delta_{F_{1}}(T)|\geq\min\bigl{\{}f(T\cup\overline{A})-|\delta_{Z}(T\cup\overline{A})|,|\delta_{G_{1}}(T)|\bigr{\}}$. Combining the inequalities, and using the definition of restriction, we obtain that $|\delta_{F_{1}}(T)|\geq\min\bigl{\{}f_{1}(T),|\delta_{G_{1}}(T)|\bigr{\}}$. Since $T$ was arbitrary, this shows that $F_{1}$ is feasible to $\mathsf{CR}\text{-}{(f_{1},G_{1})}\text{-}\mathsf{NDP}$.
A symmetric argument shows that $F_{2}$ is a feasible solution to $\mathsf{CR}\text{-}{(f_{2},G_{2})}\text{-}\mathsf{NDP}$.
Conversely, suppose that $F\supseteq\delta_{G}(A)$, $F_{1}=F(A)$ is a feasible solution to $\mathsf{CR}\text{-}{(f_{1},G_{1})}\text{-}\mathsf{NDP}$, and $F_{2}=F(\overline{A})$ is a feasible solution to $\mathsf{CR}\text{-}{(f_{2},G_{2})}\text{-}\mathsf{NDP}$. Recall that $Z=\delta_{G}(A)$. Consider any $\emptyset\neq T\subsetneq V$. Lemma 4.5 (b) can be used to readily argue that if $T\subseteq A$ or $T\subseteq\overline{A}$, then the constraint for set $T$ in $\mathsf{CR}\text{-}{(f,G)}\text{-}\mathsf{NDP}$ is satisfied. We show this for $T\subseteq A$; a symmetric argument applies for $T\subseteq\overline{A}$. If $T$ is a small-$(f_{1},G_{1})$-cut, then $F_{1}\supseteq\delta_{G_{1}}(T)$. Since $F\supseteq Z$, this implies that $F\supseteq\delta_{G}(T)$. Otherwise, we have $\operatorname{def}_{f_{1},F_{1}}(T)\leq 0$, so by Lemma 4.5 (b), we have that $\operatorname{def}_{f,F}(T)\leq\operatorname{def}_{f_{1},F_{1}}(T)\leq 0$. Thus, we always have $|\delta_{F}(T)|\geq\min\bigl{\{}f(T),|\delta_{G}(T)|\bigr{\}}$.
Now consider $T$ such that $T\cap A,T\cap\overline{A}\neq\emptyset$. We exploit the weak supermodularity of $\operatorname{def}_{f,F}$, and that $A$ has maximum deficiency, to bound $\operatorname{def}_{f,F}(T)$ in terms of the deficiency of sets that do not cross $A$, and thus show that the constraint for $T$ is satisfied. We have
$\begin{split}\operatorname{def}_{f,F}(T)+\operatorname{def}_{f,F}(A)&\leq\operatorname{def}_{f,F}(T\cap A)+\operatorname{def}_{f,F}(T\cup A)=\operatorname{def}_{f,F}(T\cap A)+\operatorname{def}_{f,F}(\overline{A}-T)\ \ \text{or}\\ \operatorname{def}_{f,F}(T)+\operatorname{def}_{f,F}(A)&\leq\operatorname{def}_{f,F}(A-T)+\operatorname{def}_{f,F}(T-A)\end{split}\vspace*{-2ex}$
where in the first inequality we also use the fact that $\operatorname{def}_{f,F}$ is symmetric. In both cases, we have $\operatorname{def}_{f,F}(T)+\operatorname{def}_{f,F}(A)\leq\operatorname{def}_{f,F}(X)+\operatorname{def}_{f,F}(Y)$ where $X\subseteq A$, $Y\subseteq\overline{A}$ and $\delta_{G}(T)\subseteq\delta_{G}(X)\cup\delta_{G}(Y)\cup Z$. As shown previously, for $S\in\{X,Y\}$, we have $|\delta_{F}(S)|\geq\min\bigl{\{}f(S),|\delta_{G}(S)|\bigr{\}}$, or equivalently, $\operatorname{def}_{f,F}(S)\leq\max\bigl{\{}0,\operatorname{def}_{f,G}(S)\bigr{\}}$. Note that $\operatorname{def}_{f,F}(A)=\operatorname{def}_{f,G}(A)>0$ since $F\supseteq\delta_{G}(A)$. Since $A$ is a maximum-deficiency small cut, this also implies that $\operatorname{def}_{f,F}(A)\geq\operatorname{def}_{f,F}(S)$ for $S\in\{X,Y\}$: we have $\operatorname{def}_{f,F}(S)\leq\max\bigl{\{}0,\operatorname{def}_{f,G}(S)\bigr{\}}\leq\max\bigl{\{}0,\operatorname{def}_{f,G}(A)\bigr{\}}=\operatorname{def}_{f,F}(A)$.
If both $X$ and $Y$ are small-$(f,G)$-cuts, then we are done since then we have that $F\supseteq\delta_{G}(X)\cup\delta_{G}(Y)\cup Z\supseteq\delta_{G}(T)$. Otherwise, since $\operatorname{def}_{f,F}(A)\geq\max\bigl{\{}\operatorname{def}_{f,F}(X),\operatorname{def}_{f,F}(Y)\bigr{\}}$, we have $\operatorname{def}_{f,F}(T)\leq\min\bigl{\{}\operatorname{def}_{f,F}(X),\operatorname{def}_{f,F}(Y)\bigr{\}}\leq 0$, since at least one of $X,Y$ is not a small-$(f,G)$-cut.
Proof 4.9 (Proof of Theorem 4.1).
Parts (a) and (b) follow easily by induction on the number of nodes, using parts (a) and (b) of Theorem 4.7. Consider part (a). The base case is when there is no small-$(f,G)$-cut, in which case, we have $V_{1}=V$, and $f_{1}=f$. Otherwise, let $A=\operatorname{argmax}_{S\subseteq V}\operatorname{def}_{f,G}(S)$. Since $f_{A}$ and $f_{\overline{A}}$ are weakly-supermodular, we can recurse and induct on $\mathsf{CR}\text{-}{(f_{A},G[{A}])}\text{-}\mathsf{NDP}$ and $\mathsf{CR}\text{-}{(f_{\overline{A}},G[{\overline{A}}])}\text{-}\mathsf{NDP}$ instances, both of which have fewer nodes, to obtain partitions of $A$ and $\overline{A}$; we combine these to obtain $V_{1},\ldots,V_{k}$. Theorem 4.7 (a) and the induction hypothesis then readily yield part (a) here. The only thing to note is that since the induction hypothesis yields $F\supseteq\delta_{G[{A}]}(V_{i})$ for a part $V_{i}\subseteq A$, and $F\supseteq\delta_{G}(A)$ by Theorem 4.7 (a), we have $F\supseteq\delta_{G}(V_{i})$. Similarly, $F\supseteq\delta_{G}(V_{i})$ for any part $V_{i}\subseteq\overline{A}$.
Part (b) follows similarly from Theorem 4.7 (b).
Part (c) follows from part (b), because it is not hard to see, using part (b), that any convex combination of feasible solutions to $(\text{{CRLP}}_{f,G})$ yields a convex combination of feasible $(\text{{NDP}}_{f_{i},G[{V_{i}}]})$-solutions, and vice versa.
Suppose $x$ is an extreme-point solution to $(\text{{CRLP}}_{f,G})$. By feasibility of $x$ and part (b), we have $x_{e}=1$ for all $e\in\delta_{G}(V_{j})$, and $x^{V_{j}}$ is feasible to $(\text{{NDP}}_{f_{j},G[{V_{j}}]})$ for all $j\in[k]$. Suppose that $x^{V_{i}}$ is not an extreme-point solution to $(\text{{NDP}}_{f_{i},G[{V_{i}}]})$ for some $i\in[k]$. Then, we can write $x^{V_{i}}$ as a convex combination $\lambda y^{1}+(1-\lambda)y^{2}$ of two solutions $y^{1}\neq x^{V_{i}}$, $y^{2}\neq x^{V_{i}}$ that are feasible to $(\text{{NDP}}_{f_{i},G[{V_{i}}]})$, where $\lambda\in[0,1]$. But then we can write $x$ as a convex combination of two feasible solutions to $(\text{{CRLP}}_{f,G})$, as follows. Define $\overline{x},\widehat{x}\in\mathbb{R}^{E}$, where $\overline{x}_{e}=\widehat{x}_{e}=x_{e}$ for all $e\in E-E(V_{i})$, and $\overline{x}_{e}=y^{1}_{e}$,  $\widehat{x}_{e}=y^{2}_{e}$ for all $e\in E(V_{i})$. So we have $\overline{x}_{e}=\widehat{x}_{e}=1$ for all $e\in\delta_{G}(V_{j})$ for all $j\in[k]$, $\overline{x}^{V_{j}}=\widehat{x}^{V_{j}}$ is feasible to $(\text{{NDP}}_{f_{j},G[{V_{j}}]})$ for all $j\in[k],j\neq i$; and $\overline{x}^{V_{i}}=y^{1}$, $\widehat{x}^{V_{i}}=y^{2}$ are feasible to $(\text{{NDP}}_{f_{i},G[{V_{i}}]})$. Again, by part (b), this yields that $\overline{x}$ and $\widehat{x}$ are feasible to $(\text{{CRLP}}_{f,G})$. But then we have $x=\lambda\overline{x}+(1-\lambda)\widehat{x}$ for feasible solutions $\overline{x}\neq x$, $\widehat{x}\neq x$, which contradicts that $x$ is an extreme-point solution to $(\text{{CRLP}}_{f,G})$.
Conversely, suppose that $x_{e}=1$ for all $e\in\delta_{G}(V_{j})$, and $x^{V_{j}}$ is an extreme-point solution to $(\text{{NDP}}_{f_{j},G[{V_{j}}]})$, for all $j\in[k]$. Then, $x$ is feasible to $(\text{{CRLP}}_{f,G})$ by part (b). Suppose $x$ is not an extreme-point solution to $(\text{{CRLP}}_{f,G})$, and $x=\lambda\overline{x}+(1-\lambda)\widehat{x}$ for feasible solutions $\overline{x}\neq x$, $\widehat{x}\neq x$. Then, again by part (b), we have $\overline{x}_{e}=\widehat{x}_{e}=1$ for all $e\in\delta_{G}(V_{j})$ for all $j\in[k]$, and so $\overline{x}^{V_{i}}\neq\widehat{x}^{V_{i}}$ for some $i\in[k]$. Also, $\overline{x}^{V_{i}}$ and $\widehat{x}^{V_{i}}$ are feasible to $(\text{{NDP}}_{f_{i},G[{V_{i}}]})$ and $x^{V_{i}}=\lambda\overline{x}^{V_{i}}+(1-\lambda)\widehat{x}^{V_{i}}$, contradicting that $x^{V_{i}}$ is an extreme-point solution to $(\text{{NDP}}_{f_{i},G[{V_{i}}]})$.
One may wonder if the ${(f_{i},G[{V_{i}}])}\text{-}\mathsf{NDP}$-instances in the decomposition of Theorem 4.1 can be specified more directly or succinctly. We prove some properties of the decomposition given by its proof in Section 6, which as a consequence, enable one to do so (see Lemma 6.6). But we need our splitting approach to argue why this succinct description fulfills the properties in Theorem 4.1. We discuss this, and other subtleties in Section 6.1. The properties we establish also imply that the decomposition in Theorem 4.1 can be computed efficiently given suitable algorithmic access to the base cut-requirement function $f$.
5 Algorithmic implication: LP-relative $2$-approximation algorithm
We now exploit the decomposition result stated in Theorem 4.1, and Corollary 4.2, to obtain a $2$-approximation algorithm. We assume that we have a separation oracle for $(\text{{CRLP}}_{f,G})$, which implies that we can efficiently find an extreme-point optimal solution to $(\text{{CRLP}}_{f,G})$ and any LP that arises in a subsequent iteration after fixing some $x_{e}$ variables to $1$ (Fact 7).
 Algorithm CRNDP-Alg. The input is $\bigl{(}G=(V,E),\{c_{e}\geq 0\},f:2^{V}\mapsto\mathbb{Z}\bigr{)}$, where $f$ is weakly-supermodular (and symmetric, normalized). We assume there is a separation oracle for $(\text{{CRLP}}_{f,G})$.
  1. A1.
    Initialize $F\leftarrow\emptyset$, $G^{\prime}=(V,E^{\prime})\leftarrow G$, and $f^{\prime}\leftarrow f$.
  2. A2.
    While $F$ is not a feasible solution (which can be detected using the separation oracle), repeat:
    1. A2.1
      Find an extreme-point optimal solution $\widehat{x}$ to $(\text{{CRLP}}_{f^{\prime},G^{\prime}})$.
    2. A2.2
      Let $Z_{1}=\{e\in E^{\prime}:\widehat{x}_{e}\geq 1/2\}$. Update $F\leftarrow F\cup Z_{1}$, $E^{\prime}\leftarrow E^{\prime}-Z_{1}$, and $f^{\prime}(S)=f^{\prime}(S)-|\delta_{Z_{1}}(S)|$ for all $S\subseteq V$.
  3. A3.
    Return $F$.
Remark 5.1.
The collection of small cuts does not change across iterations; that is, $S$ is a small-$(f,G)$-cut iff it is a small-$(f^{\prime},G^{\prime})$-cut in some other iteration. This is because, for any $S\subseteq V$, we change $|\delta_{G^{\prime}}(S)|$ and $f^{\prime}(S)$ by the same amount in every iteration. In the first iteration, we have $\widehat{x}_{e}=1$ for every $e\in\delta_{G}(S)$ and every small cut $S$, and so $\delta_{G^{\prime}}(S)=\emptyset$ in any subsequent iteration. So we do not have any constraints for small cuts in subsequent iterations, and this causes problems with uncrossing tight constraints. Nevertheless, we are able to argue, due to our decomposition result, that there is always some edge with $\widehat{x}_{e}\geq 1/2$.
Theorem 5.2.
[…]
Proof 5.3.
[…]
Since $(\text{{CRLP}}_{f^{\mathsf{SNDP}},G})$ admits a polytime separation oracle (Lemma 2.3), we obtain the following result as a corollary.
Theorem 5.4.
There is a $2$-approximation algorithm for $\mathsf{CR}\text{-}\mathsf{SNDP}$.
6 Properties of the decomposition in Theorem 4.1
Throughout this section, $f$ and $G$ always denote the base cut-requirement function $f$ and graph $G$ that are given as input to $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$. Theorem 4.1 does not pinpoint a specific decomposition, and only mentions some properties that the $(f_{i},V_{i})$ tuples should satisfy. By “the decomposition in Theorem 4.1,” we mean the decomposition given by its proof, wherein we obtain a maximum deficiency set $A\subseteq V$ as in Theorem 4.7, and recurse on the $\mathsf{CR}\text{-}{(f_{A},G[{A}])}\text{-}\mathsf{NDP}$ and $\mathsf{CR}\text{-}{(f_{\overline{A}},G[{\overline{A}}])}\text{-}\mathsf{NDP}$-instances. This process can be conveniently described in terms of a decomposition tree as follows. Each node of the tree is labeled by a tuple $(h,S)$, where $S\subseteq V$ and $g:2^{S}\mapsto\mathbb{Z}$ is a weakly-supermodular, symmetric, normalized cut-requirement function on $S$; we will often simply say node $(h,S)$. The root of the tree is labeled $(f,V)$. For each (current) leaf node $(h,S)$, we find $A\subseteq S$ such that $\operatorname{def}_{h,G[{S}]}(A)=\max_{T\subseteq S}\operatorname{def}_{h,G[{S}]}(T)$. If $\operatorname{def}_{f,G[{S}]}(A)\leq 0$, then there is no small-$(h,G[{S}])$-cut and $(h,S)$ gets fixed as a leaf node. Otherwise, we create two children of $(h,S)$ labeled $(h_{1},A)$ and $(h_{2},S-A)$, where $h_{1}$ and $h_{2}$ are the restrictions of $h$ to $A$ and $S-A$ respectively (see Definition 4.4), and continue. We call this tree the canonical decomposition tree, and refer to the decomposition given by its leaves as the canonical decomposition.
We establish some useful properties of the canonical decomposition tree that may be of independent interest. In particular, one can infer from these properties that the canonical decomposition tree can be computed efficiently given suitable algorithmic access to $f$. This suggests another way of obtaining an algorithmic result for $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$, namely, by computing the canonical decomposition efficiently and then (approximately) solving the ${(f_{i},G[{V_{i}}])}\text{-}\mathsf{NDP}$-instances. In particular, if the resulting $f_{i}$ functions have some additional structure, then these ${(f_{i},G[{V_{i}}])}\text{-}\mathsf{NDP}$-instances might admit other, possibly more-efficient approximation algorithms (e.g., via the primal-dual method), which would yield an alternate (to iterative rounding) way of obtaining an approximation result for $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$.
Another consequence of the properties we establish is that they allow us to succinctly describe the $(f_{i},V_{i})$ tuples given by the leaves of the canonical decomposition tree (see Lemmas 6.6 and 6.8). While this yields a self-contained, succinct description of the canonical decomposition, we emphasize that in order to prove that the ${f}\text{-}\mathsf{NDP}$-instances given by this succinct description satisfy the properties stated in Theorem 4.1, we need to go via the canonical decomposition and exploit Lemma 6.1 (and thus our approach of iteratively simplifying the instance via splitting along maximum-deficiency small cuts). So having this succinct description does not simplify the underlying arguments; see also Section 6.1.1.
Lemma 6.1.
Let $(h,S)$ be a node of the canonical decomposition tree. Let the nodes on the path from $(h,S)$ to the root be labeled $(h_{0},S_{0})=(h,S),(h_{1},S_{1}),\ldots,(h_{\ell},S_{\ell})=(f,V)$. Define $R_{I}:=\bigcup_{i\in I}(S_{i}-S_{i-1})$. Note that $R_{\emptyset}:=\emptyset$. For $i\in[\ell]$, we abbreviate $R_{\{i\}}$ to $R_{i}$, so $R_{i}=S_{i}-S_{i-1}$.
  1. (a)
    For any $\emptyset\neq T\subsetneq S$, we have $\operatorname{def}_{h,G[{S}]}(T)=\max\,\bigl{\{}\operatorname{def}_{f,G}(T\cup R_{I}):I\subseteq[\ell]\bigr{\}}$.
  2. (b)
    For any $i,j\in\{0\}\cup[\ell]$ with $i\leq j$, and any $T\subseteq S_{i}$, we have $\operatorname{def}_{h_{j},G[{S_{j}}]}(T)\leq\operatorname{def}_{h_{i},G[{S_{i}}]}(T)$.
  3. (c)
    Let $s,t\in S$. We have
    $\max\,\bigl{\{}\operatorname{def}_{h,G[{S}]}(T):\ T\subseteq S\text{ is an $s$-$t$ cut}\bigr{\}}=\max\,\bigl{\{}\operatorname{def}_{f,G}(T):\ T\subseteq V\text{ is an $s$-$t$ cut}\bigr{\}}.$
    Moreover, if an $s$-$t$ cut $T\subseteq V$ attains the maximum on the right, then $T\cap S$ attains the maximum on the left.
Proof 6.2.
We prove part (a) by induction on $\ell$, which we call the depth of $(h,S)$. The base case when $\ell=0$ is trivially true. Suppose the statement holds whenever the depth is at most $d$, and consider a node $(h,S)$ with depth $\ell=d+1$. Then $h$ is the restriction of $h_{1}$ to $S$, so letting $Z=\delta_{G[{S_{1}}]}(S)$, we have
$\operatorname{def}_{h,G[{S}]}(T)$ $=\max\bigl{\{}h_{1}(T)-|\delta_{Z}(T)|,h_{1}(T\cup S_{1}-S)-|\delta_{Z}(T\cup S_{1}-S)|\bigr{\}}-|\delta_{G[{S}]}(T)|$
$=\max\bigl{\{}\operatorname{def}_{h_{1},G[{S_{1}}]}(T),\operatorname{def}_{h_{1},G[{S_{1}}]}(T\cup R_{1})\bigr{\}}.$ (7)
By the induction hypothesis, we have
$\begin{split}\operatorname{def}_{h_{1},G[{S_{1}}]}(T)&=\max\,\bigl{\{}\operatorname{def}_{f,G}(T\cup R_{I}):I\subseteq\{2,\ldots,\ell\}\bigr{\}},\qquad\text{and}\\ \operatorname{def}_{h_{1},G[{S_{1}}]}(T\cup R_{1})&=\max\,\bigl{\{}\operatorname{def}_{f,G}(T\cup R_{1}\cup R_{I}):I\subseteq\{2,\ldots,\ell\}\bigr{\}}.\end{split}$
Plugging these in (7), and noting that $R_{1}\cup R_{I}=R_{\{1\}\cup I}$ for $I\subseteq\{2,\ldots,\ell\}$ yields $\operatorname{def}_{h,G[{S}]}(T)=\max\,\bigl{\{}\operatorname{def}_{f,G}(T\cup R_{I}):I\subseteq[\ell]\bigr{\}}$, completing the induction step, and the proof.
Part (b) follows directly from part (a), since the collection of sets $\bigl{\{}T\cup R_{I}:I\subseteq\{i+1,\ldots,\ell\}\bigr{\}}$ considered in the computation of $\operatorname{def}_{h_{i},G[{S_{i}}]}(T)$ is a superset of the collection $\bigl{\{}T\cup R_{I}:I\subseteq\{j+1,\ldots,\ell\}\bigr{\}}$ considered for $\operatorname{def}_{h_{j},G[{S_{j}}]}(T)$.
We now prove part (c). Let LHS = $\max\,\bigl{\{}\operatorname{def}_{h,G[{S}]}(T):\ T\subseteq S\text{ is an $s$-$t$ cut}\bigr{\}}$ and RHS = $\max\,\bigl{\{}\operatorname{def}_{f,G}(T):\ T\subseteq V\text{ is an $s$-$t$ cut}\bigr{\}}$. By part (a), for any $s$-$t$ cut $T\subseteq S$, we can find some $s$-$t$ cut $T_{1}\subseteq V$ such that $\operatorname{def}_{h,G[{S}]}(T)=\operatorname{def}_{f,G}(T_{1})$. So LHS $\leq$ RHS.
For the other direction, fix any $s$-$t$ cut $T_{1}\subseteq V$. We show that $\operatorname{def}_{f,G}(T_{1})\leq\operatorname{def}_{h,G[{S}]}(T_{1}\cap S)$, which shows that RHS $\leq$ LHS, and also yields the desired statement relating maximizers of the RHS and the LHS.
Recall that $(h_{1},S_{1}),(h_{2},S_{2}),\ldots,(h_{\ell},S_{\ell})=(f,V)$ are the nodes on the path from $(h,S)$ to the root, with $(h_{1},S_{1})$ being the parent of $(h,S)$. Let $a\in\{0,1,\ldots,\ell\}$ be the smallest index such that $T_{1}\subseteq S_{a}$. Define $X_{i}:=T_{1}\cap S_{i}$ for all $i\in[a]$ (so $X_{a}:=T_{1}$). We exploit the weak-supermodularity (and symmetry) of the $\operatorname{def}_{h_{i},G[{S_{i}}]}(\cdot)$ functions to show that
$\operatorname{def}_{h_{i},G[{S_{i}}]}(X_{i})\leq\operatorname{def}_{h_{i-1},G[{S_{i-1}}]}(X_{i-1})\qquad\forall i\in[a].$ (8)
Clearly, (8) implies that $\operatorname{def}_{h_{a},G[{S_{a}}]}(T_{1})\leq\operatorname{def}_{h,G[{S}]}(T_{1}\cap S)$. We also have $\operatorname{def}_{f,G}(T_{1})\leq\operatorname{def}_{h_{a},G[{S_{a}}]}(T_{1})$, which finishes the proof.
We now prove (8). We have $\operatorname{def}_{h_{i},G[{S_{i}}]}(X_{i})+\operatorname{def}_{h_{i},G[{S_{i}}]}(S_{i-1})\leq\operatorname{def}_{h_{i},G[{S_{i}}]}(U_{i-1})+\operatorname{def}_{h_{i},G[{S_{i}}]}(Y_{i-1})$ where $U_{i-1}=X_{i}\cap S_{i-1}$, $Y_{i-1}=X_{i}\cup S_{i-1}$, or $U_{i-1}=S_{i-1}-X_{i}$, $Y_{i-1}=X_{i}-S_{i-1}$. By the definition of $S_{i-1}$, we have $\operatorname{def}_{h_{i},G[{S_{i}}]}(S_{i-1})\geq\operatorname{def}_{h_{i},G[{S_{i}}]}(Y)$ for all $Y\subseteq S_{i}$. So the above inequality yields that $\operatorname{def}_{h_{i},G[{S_{i}}]}(X_{i})\leq\operatorname{def}_{h_{i},G[{S_{i}}]}(U_{i-1})$. Also, by part (b), we have $\operatorname{def}_{h_{i},G[{S_{i}}]}(U_{i-1})\leq\operatorname{def}_{h_{i-1},G[{S_{i-1}}]}(U_{i-1})$. We know that $U_{i-1}$ is $X_{i}\cap S_{i-1}=X_{i-1}$ or $S_{i-1}-X_{i}$. So since $\operatorname{def}_{h_{i-1},G[{S_{i-1}}]}(\cdot)$ is symmetric, we therefore obtain that $\operatorname{def}_{h_{i-1},G[{S_{i-1}}]}(U_{i-1})=\operatorname{def}_{h_{i-1},G[{S_{i-1}}]}(X_{i}\cap S_{i-1})=\operatorname{def}_{h_{i-1},G[{S_{i-1}}]}(X_{i-1})$. Combining everything, we obtain that (8) holds for index $i$.
Lemma 6.1 (c) indicates that finding maximum-deficiency $s$-$t$ cuts is the natural algorithmic problem that one needs to solve in order to find the canonical decomposition tree, and the resulting decomposition. Formally, in the max-deficiency $s$-$t$ cut problem, we are given $s,t\in V$ and we need to return an $s$-$t$ cut $A\subseteq V$ that maximizes $\operatorname{def}_{f,G}(T)$ over all $s$-$t$ cuts $T\subseteq V$ (and its deficiency $\operatorname{def}_{f,G}(A)$). More generally, in the weighted max-deficiency $s$-$t$ cut problem, we are also given $w\in\mathbb{R}_{+}^{E}$, and we want find an $s$-$t$ cut $A\subseteq V$ that maximizes $\operatorname{wdef}_{f,G}(S):=f(T)-w\bigl{(}\delta_{G}(T)\bigr{)}$ over all $s$-$t$ cuts $T\subseteq V$.
Theorem 6.3.
Given a polytime algorithm for the max-deficiency $s$-$t$ cut problem, we can efficiently compute the canonical decomposition tree, and hence the decomposition in Theorem 4.1.
Proof 6.4.
We only need to show that given a node $(h,S)$ of the current canonical decomposition tree, how to find $A\subseteq S$ such that $\operatorname{def}_{h,G[{S}]}(A)=\max_{T\subseteq S}\operatorname{def}_{h,G[{S}]}(T)$. Fix some $s\in S$. Clearly, if we can find $\max\,\{\operatorname{def}_{h,G[{S}]}(T):T\subseteq S\text{ is an $s$-$t$ cut}\}$, then we can find $A$ by taking the maximum over these $s$-$t$ cuts. Using Lemma 6.1 (c), we can optimize over $s$-$t$ cuts contained in $S$ by finding a max $s$-$t$ deficient cut with respect to $(f,G)$ using , and taking its intersection with $S$.
Furthermore, note that if can solve weighted max $s$-$t$ deficient cut in polytime, then for any leaf $(f_{i},S)$ of the canonical decomposition tree, we can give an efficient separation oracle for the LP $(\text{{NDP}}_{f_{i},G[{S}]})$. This is because exactly as in Lemma 6.1 (c), we have $\max\,\bigl{\{}\operatorname{wdef}_{f_{i},G[{S}]}(T):\ T\subseteq S\text{ is an $s$-$t$ cut}\bigr{\}}=\max\,\bigl{\{}\operatorname{wdef}_{f,G}(T):\ T\subseteq V\text{ is an $s$-$t$ cut}\bigr{\}}$, and if $T$ attains the maximum on the right, then $T\cap S$ attains the maximum on the left. So we can find $\max_{T\subseteq S}\bigl{(}f_{i}(T)-x\bigl{(}\delta_{G[{S}]}(T)\bigr{)}$ by performing $|S|$ weighted max-$s$-$t$-deficient-cut computations (with respect to $(f,G)$).
For $\mathsf{CR}\text{-}\mathsf{SNDP}$, where the base cut-requirement function is $f^{\mathsf{SNDP}}$, finding a weighted $s$-$t$ deficient cut amounts to a suitable min-cut problem. Recall that the $\mathsf{SNDP}$ input specifies $k$ tuples $(s_{i},t_{i},r_{i})$, where $s_{i},t_{i}\in V$, and $r_{i}\in\mathbb{Z}_{+}$ for all $i$, and $f^{\mathsf{SNDP}}(S)=\max\,\{r_{i}:|S\cap\{s_{i},t_{i}\}|=1\}$. So to solve $\max\,\{f^{\mathsf{SNDP}}(T)-w\bigl{(}\delta_{G}(T)\bigr{)}:\ T\subseteq V\text{ is an $s$-$t$ cut}\bigr{\}}$, we consider each $i\in[k]$, and find $T_{i}$ that minimizes $w\bigl{(}\delta_{G}(T)\bigr{)}$ over all sets $T\subseteq V$ such that $|T\cap\{s_{i},t_{i}\}|=1=|T\cap\{s,t\}|$, by solving a min-cut problem. We then return $\max_{i\in[k]}\bigl{(}r_{i}-w\bigl{(}\delta_{G}(T_{i})\bigr{)}$. We therefore obtain the following corollary.
Corollary 6.5.
For $\mathsf{CR}\text{-}\mathsf{SNDP}$, we can find the canonical decomposition tree in polytime. Also, for any leaf $(f_{i},S)$ of this tree, we have a polytime separation oracle for $(\text{{NDP}}_{f_{i},G[{S}]})$.
6.1 Succinct description of the canonical decomposition
As noted earlier, one structural consequence of Lemma 6.1 is that one can give a succinct description of the $(f_{i},V_{i})$ tuples obtained at the leaves of the canonical decomposition tree. We show that the $V_{1},\ldots,V_{k}$ leaf-set partition is precisely the partition given by the (node-sets of the) components of $G-Z$, where $Z:=\bigcup_{S\subseteq V:\operatorname{def}_{f,G}(S)>0}\delta_{G}(S)$ is the set of all edges on the boundary of some small cut. We can also succinctly specify the $f_{i}:2^{V_{i}}\mapsto\mathbb{Z}$ functions for $i\in[k]$ (Lemma 6.8).
Lemma 6.6.
The leaf sets $V_{1},\ldots,V_{k}$ of the canonical decomposition correspond to the node-sets of the components of $G-Z$.
Proof 6.7.
We first argue that the components of $G-Z$ yield a refinement of the $V_{1},\ldots,V_{k}$ partition. To do so, consider any node $(h,S)\neq(f,V)$ of the canonical decomposition tree, and let $(h_{1},S_{1})$ be the parent of $(h,S)$. We claim that $\delta_{G[{S_{1}}]}(S)\subseteq Z$. Applying this to every node of the canonical decomposition tree yields that $\delta_{G}(V_{i})\subseteq Z$ for every leaf-set $V_{i}$. To see the claim, applying part (a) of Lemma 6.1 to $(h_{1},S_{1})$ taking $T=S$, since $\operatorname{def}_{h_{1},G[{S_{1}}]}(S)>0$, we obtain that there is some small-$(f,G)$-cut $A\subseteq V$ such that $S=A\cap S_{1}$. Thus, $\delta_{G[{S_{1}}]}(S)\subseteq\delta_{G}(A)\subseteq Z$.
We next argue that no edge of $Z$ is internal to a leaf set, which will finish the proof. Suppose not, and there is some $e=uv\in Z$ with both $u,v\in V_{i}$ for some $i\in[k]$. Then, by the definition of $Z$, there is some small-$(f,G)$-cut $T$ such $e\in\delta_{G}(T)$. But then taking $s=u,t=v$, and by Lemma 6.1 (c), there is some small-$(f_{i},G[{V_{i}}])$-cut $T^{\prime}\subsetneq V_{i}$ separating $s$ and $t$. This yields a contradiction since $(f_{i},V_{i})$ is a leaf of the canonical decomposition tree.
We point out that for a leaf-set $V_{i}$, while $\delta_{G}(V_{i})\subseteq Z$, it need not be that $V_{i}$ is a small-$(f,G)$-cut. For instance, if the base cut-requirement function $f$ models an $s$-$t$ connectivity problem, then every small-$(f,G)$-cut must be an $s$-$t$ cut, and so unless there are only two leaf sets, this will not hold.
Lemma 6.8.
Let $(f_{1},V_{1}),\ldots,(f_{k},V_{k})$ be the leaves of the canonical decomposition tree. Consider $i\in[k]$, and let $\mathcal{C}_{i}:=\bigl{\{}\bigcup_{j\in I}V_{j}:I\subseteq[k]-\{i\}\bigr{\}}$. Define $\gamma_{i}:2^{V_{i}}\mapsto\mathbb{Z}$ as follows:
$\gamma_{i}(\emptyset)=\gamma_{i}(V_{i}):=0,\qquad\gamma_{i}(S):=\max\,\bigl{\{}f(S\cup X)-|\delta_{Z}(S\cup X)|:X\in\mathcal{C}_{i}\bigr{\}}\quad\forall\emptyset\neq S\subsetneq V_{i}.$
Then $\gamma_{i}$ coincides with $f_{i}$. the function obtained from the canonical decomposition, for all $i\in[k]$.
Proof 6.9.
Let $(h_{1},S_{1}),\ldots,(h_{\ell},S_{\ell})=(f,V)$ be the nodes of the canonical decomposition tree on the path from $(f_{i},V_{i})$ to the root, with $(h_{1},S_{1})$ being the parent of $(f_{i},V_{i})$. Consider any $\emptyset\neq T\subsetneq V_{i}$. We first show that $\gamma_{i}(T)\geq f_{i}(T)$, or equivalently $\operatorname{def}_{\gamma_{i},G[{V_{i}}]}(T)\geq\operatorname{def}_{f_{i},G[{V_{i}}]}(T)$. This is because, by Lemma 6.1 (a), one can infer that $f_{i}$ is defined similarly to $\gamma_{i}$, except that we restrict the sets $X\in\mathcal{C}_{i}$ over which the maximum is taken. Recall that for $I\subseteq[\ell]$, we define $R_{I}:=\bigcup_{j\in I}(S_{j}-S_{j-1})$, and $R_{j}$ denotes $R_{\{j\}}:=S_{j}-S_{j-1}$, for $j\in[\ell]$. Observe that for any $j\in[\ell]$, we have that $S_{j}-S_{j-1}$ is a union of some leaf-sets other than $V_{i}$. Therefore, $S_{j}-S_{j-1}\in\mathcal{C}_{i}$ for every $j\in[\ell]$, and hence $R_{I}\in\mathcal{C}_{i}$ for all $I\subseteq[\ell]$. We have
$\operatorname{def}_{f_{i},G[{V_{i}}]}(T)=\max\,\bigl{\{}\operatorname{def}_{f,G}(T\cup R_{I}):I\subseteq[\ell]\bigr{\}}\leq\max\,\bigl{\{}\operatorname{def}_{f,G}(T\cup X):X\in\mathcal{C}_{i}\bigr{\}}=\operatorname{def}_{\gamma_{i},G[{V_{i}}]}(T).$
The first inequality is because $R_{I}\in\mathcal{C}_{i}$ for every $I\subseteq[\ell]$, and the last equation follows because $\delta_{G}(T\cup X)$ is the union of two disjoint sets, $\delta_{Z}(T\cup X)$ and $\delta_{G[{V_{i}}]}(T)$.
We now argue that $\gamma_{i}(T)\leq f_{i}(T)$. Let $\operatorname{def}_{\gamma_{i},G[{V_{i}}]}(T)=\operatorname{def}_{f,G}(T\cup X)$, where $X\in\mathcal{C}_{i}$. Let $T_{1}=T\cup X$. In the proof of part (c) of Lemma 6.1, we argued that $\operatorname{def}_{f,G}(T_{1})\leq\operatorname{def}_{f_{i},G[{V_{i}}]}(T_{1}\cap V_{i})$ (see (8)). But $T_{1}\cap V_{i}=T$, so $\operatorname{def}_{\gamma_{i},G[{V_{i}}]}(T)\leq\operatorname{def}_{f_{i},G[{V_{i}}]}(T)$.
6.1.1 Some remarks and discussion
The reader may wonder whether one can directly argue that the $(\gamma_{i},V_{i})$ tuples specified succinctly in Lemmas 6.6 and 6.8 satisfy the properties stated in Theorem 4.1. It is easy to show that if $F$ is feasible for $\mathsf{CR}\text{-}{(f,G)}\text{-}\mathsf{NDP}$, then we must have $F\supseteq\delta_{G}(V_{i})$ and that $F(V_{i})$ is a feasible solution to ${(\gamma_{i},G[{V_{i}}])}\text{-}\mathsf{NDP}$, for all $i\in[k]$. This is simply because, for any $i\in[k]$, any set $\emptyset\neq S\subsetneq V_{i}$, and any $X\in\mathcal{C}_{i}$, we have $|\delta_{F}(S\cup X)|\geq\min\bigl{\{}f(S\cup X),|\delta_{G}(S\cup X)|\bigr{\}}=f(S\cup X)$ (the last equality is because $\delta_{G}(S\cup X)\not\subseteq Z$), which implies that $|\delta_{F(V_{i})}(S)|\geq f(S\cup X)-|\delta_{Z}(S\cup X)|$.
The other direction is more tricky. Since $F\supseteq\bigcup_{i\in[k]}\delta_{G}(V_{i})$, we can focus on sets that are not a union of leaf-sets, and hence are not small-$(f,G)$-cuts. Recall that we say that two sets $A,B$ cross if $A-B,A\cap B$ and $B-A$ are all non-empty. Suppose $S\subseteq V$ crosses some $r$ leaf sets. Similar to the proof of Lemma 4.7, one potential approach would be to utilize the weak supermodularity of $\operatorname{def}_{f,F}$ to move to sets that cross fewer leaf sets, and thereby use induction on $r$ to establish that $\operatorname{def}_{f,F}(S)\leq 0$. There are some subtle issues that arise here that we want to highlight.
The base cases, $r=0$ or $r=1$ are easy enough to argue, using the fact that $F(V_{i})$ is a feasible ${(\gamma_{i},G[{V_{i}}])}\text{-}\mathsf{NDP}$-solution for all $i\in[k]$. So suppose $r\geq 2$, A natural idea would be to pick some leaf set that $S$ crosses, say $V_{i}$, and apply weak supermodularity to $S$ and $V_{i}$, since we know that all four sets $S\cap V_{i},S\cup V_{i},S-V_{i},V_{i}-S$ cross fewer than $r$ leaf sets, and none of these sets is a union of leaf-sets and hence a small-$(f,G)$-cut. This would yield, for $\{X,Y\}=\{S\cap V_{i},S\cup V_{i}\}$ or $\{X,Y\}=\{S-V_{i},V_{i}-S\}$, that $\operatorname{def}_{f,F}(S)+\operatorname{def}_{f,F}(V_{i})\leq\operatorname{def}_{f,F}(X)+\operatorname{def}_{f,F}(Y)\leq 0$. However, since $V_{i}$ need not be a small-$(f,G)$-cut, this does not let us infer anything about $\operatorname{def}_{f,F}(S)$. To fix this, one could try to pick some small-$(f,G)$-cut $T\supseteq V_{i}$. But then we need not have that all the sets $S\cap T,S\cup T,S-T,T-S$ cross fewer than $r$ leaf sets. In particular, if $T$ contains all the same leaf-sets that $S$ crosses, then $S\cap T$ and $T-S$ would also cross all of these leaf sets.
The key missing ingredient needed to make this approach work is the following leaf-set separation property: for any two leaf sets, there is some small-$(f,G)$-cut that separates them. This can be shown using insights from Lemma 6.1; we do not know of any other way of showing this. Given this property, one can complete the induction proof as follows. Let $V_{i}$ and $V_{j}$ be two leaf sets that $S$ crosses. Now let $T$ be a small-$(f,G)$-cut separating $V_{i}$ and $V_{j}$. Now, it is the case that all four sets $S\cap T,S\cup T,S-T,T-S$ cross fewer than $r$ leaf sets, so by weak-supermodularity and the induction hypothesis, we obtain that $\operatorname{def}_{f,F}(S)+\operatorname{def}_{f,F}(T)\leq 0$, which implies that $\operatorname{def}_{f,F}(S)\leq 0$, since $T$ is a small-$(f,G)$-cut.
To prove the leaf-set separation property, let $(f_{i},V_{i})$ and $(f_{j},V_{j})$ be two leaves of the canonical decomposition tree. Let $(h,S)$ be their least common ancestor, and let $(h_{1},A)$, $(h_{2},S-A)$ be the children of $(h,S)$. Clearly, $A$ separates $V_{i}$ and $V_{j}$. We have that $0<\operatorname{def}_{h,G[{S}]}(A)=\max\,\{\operatorname{def}_{h,G[{S}]}(T):T\subseteq S\}$. So for any $s\in A$, $t\in S-A$, by Lemma 6.1 (c), we may assume that $A=T\cap S$, where $\operatorname{def}_{f,G}(T)=\operatorname{def}_{h,G[{S}]}(A)$. It follows that $T$ is a small-$(f,G)$-cut that separates $V_{i}$ and $V_{j}$.
7 Hardness result
We now show that cut-relative $\mathsf{SNDP}$ and path-relative $\mathsf{SNDP}$ are APX-hard, even when the base $\mathsf{SNDP}$ instance involves only one $s$-$t$ pair, by showing that the $k$-edge connected subgraph ($k\text{-}\mathsf{ECSS}$) problem can be cast as a special case of these problems. (This hints at the surprising amount of modeling power of {cut, path}- relative $\mathsf{SNDP}$.) Previously, even NP-hardness of path-relative $\mathsf{SNDP}$ in the $s$-$t$ case was not known.
Let $\bigl{(}G=(V,E),\{c_{e}\}_{e\in E}\bigr{)}$ be an instance of $k\text{-}\mathsf{ECSS}$. We may assume that $G$ is at least $k$-edge connected, i.e., $|\delta_{G}(S)|\geq k$ for all $\emptyset\neq S\subsetneq V$, as otherwise there is no feasible solution (and this can be efficiently detected).
Consider the following path-relative $\mathsf{SNDP}$ instance. We add a source $s$ and sink $t$, and edges $sv,vt$ of cost 00, for all $v\in V$. Let $G^{\prime}=(V^{\prime},E^{\prime})$ denote this graph. The base-$\mathsf{SNDP}$ instance asks for $(k+n)$-edge connectivity between $s$ and $t$, where $n=|V|$. That is, the base cut-requirement function $f$ is given by $f(S^{\prime})=k+n$ if $S^{\prime}\subseteq V^{\prime}$ is an $s$-$t$ cut, and $f(S^{\prime})=0$ otherwise. We show that path-relative $\mathsf{SNDP}$ and cut-relative $\mathsf{SNDP}$ are actually equivalent in this case, and that they encode $k\text{-}\mathsf{ECSS}$ on the graph $G$.
Theorem 7.1.
$H\subseteq E^{\prime}$ is a feasible solution to the above {cut, path}-relative $\mathsf{SNDP}$ instance $\iff$ $H\supseteq\delta_{G^{\prime}}(s)\cup\delta_{G^{\prime}}(t)$ and $H(V)$ is a feasible $k\text{-}\mathsf{ECSS}$-solution for $G$. Hence, this {cut,path}-relative $\mathsf{SNDP}$ instance is the same as $k\text{-}\mathsf{ECSS}$ on the graph $G$.
Proof 7.2.
Recall that we are assuming that $G$ is (at least) $k$-edge connected. The following observation will be handy. The only small cuts in $G^{\prime}$ are $\{s\}$ and $\{t\}$ (and their complements). For any other $s$-$t$ cut $S^{\prime}\subseteq V^{\prime}$, taking $S=S^{\prime}\cap V$, we have that $\emptyset\neq S\subsetneq V$. So we have $|\delta_{G^{\prime}}(S^{\prime})|=|\delta_{G}(S)|+|S|+|V-S|\geq k+n$.
$\Leftarrow$ direction. We argue that $H$ is feasible for cut-relative $\mathsf{SNDP}$, which also implies that it is feasible for path-relative $\mathsf{SNDP}$. Clearly, $H$ covers both small cuts. For any other $s$-$t$ cut $S^{\prime}\subseteq V^{\prime}$, taking $S=S^{\prime}\cap V$, we have $|\delta_{H}(S^{\prime})|\geq|\delta_{H(V)}(S)|+|S|+|V-S|\geq k+n$, since $H(V)$ is a feasible $k\text{-}\mathsf{ECSS}$-solution for $G$.
$\Rightarrow$ direction. Suppose that $H$ is feasible for path-relative $\mathsf{SNDP}$. If there is some edge $sv\notin H$, then consider the fault-set $F=\delta_{G^{\prime}}(s)-\{sv\}$. Then $H-F$ clearly has no $s$-$t$ path, since $\delta_{H-F}(s)=\emptyset$. But $G^{\prime}-F$ has an $s$-$t$ path $sv,vt$. So we must have $H\supseteq\delta_{G^{\prime}}(s)$. A similar argument shows that $H\supseteq\delta_{G^{\prime}}(t)$.
We next argue that $|\delta_{H}(S^{\prime})|\geq n+k$ for every $s$-$t$ cut $S^{\prime}$ for which $\emptyset\neq S:=S^{\prime}\cap V\subsetneq V$. This is equivalent to showing that $|\delta_{H(V)}(S)|\geq k$, since $|\delta_{H}(S^{\prime})\cap\delta_{G^{\prime}}(s)|+|\delta_{H}(S^{\prime})\cap\delta_{G^{\prime}}(t)|=n$. Assume that $s\in S^{\prime}$ without loss of generality. Suppose $|\delta_{H(V)}(S)|<k$. Since $|\delta_{G}(S)|\geq k$, there is some edge $uv\in\delta_{G}(S)-\delta_{H(V)}(S)$. Suppose $u\in S$. Consider the fault-set $F=\delta_{H(V)}(S)\cup\{sw:w\in V-S\}\cup\{wt:w\in S\}$. Then $|F|<k+n$ and $H-F$ has no $s$-$t$ path, since we have constructed $F$ to ensure that $\delta_{H-F}(S^{\prime})=\emptyset$. However, $G-F$ has an $s$-$t$ path: $su,uv,vt$. This contradicts that $H$ is feasible for path-relative $\mathsf{SNDP}$.
So we have shown both that $H(V)$ is feasible $k\text{-}\mathsf{ECSS}$ solution for $G$, and that $H$ is feasible for cut-relative $\mathsf{SNDP}$.
Since the cost of the edges in $E^{\prime}-E$ is $0$, the above {cut, path}-relative $\mathsf{SNDP}$ instance is exactly the same as $k\text{-}\mathsf{ECSS}$ on the graph $G$.
Appendix A Non-equivalence of path-relative $\mathsf{SNDP}$ and cut-relative $\mathsf{SNDP}$
We consider the same graph and base-SNDP problem as in Fig. 1. The graph is reproduced below for convenience, and recall that the base cut-requirement function models $s$-$t$ $2$-edge connectivity: so $f(S)=2$ if $S$ is an $s$-$t$ cut, and $f(S)=0$ otherwise.
Figure 2: An instance where path-relative $\mathsf{SNDP}$ and cut-relative $\mathsf{SNDP}$ are not equivalent. The number labeling an edge gives the cost of the edge.
A feasible solution to path-relative $\mathsf{SNDP}$ is given by the edge-set $H_{1}=\{su,uw,sw,wt\}$, which has cost $3$. (This is in fact an optimal solution to path-relative $\mathsf{SNDP}$.)
However, $H_{1}$ is not feasible for cut-relative $\mathsf{SNDP}$. This is because, we require $\delta_{H}(\{s,u,w\})\geq\min\{2,3\}=2$ in any feasible $\mathsf{CR}\text{-}\mathsf{SNDP}$ solution $H$. Moreover, one can infer that any feasible $\mathsf{CR}\text{-}\mathsf{SNDP}$-solution must include all but one of the edges from $\{su,uw,sw,sv,vw\}$. This is because for any $F\subseteq\{su,uw,sw,sv,vw\}$ with $|F|=2$, one can verify that there is an $s$-$t$ cut $S$ such that $|\delta_{G}(S)|=3$ and $F\subseteq\delta_{G}(S)$; so $|\delta_{G-F}(S)|\leq 1$, which implies that there is no feasible solution that excludes $F$. A feasible $\mathsf{CR}\text{-}\mathsf{SNDP}$-solution must also include the edge $wt$, so the optimal value for cut-relative $\mathsf{SNDP}$ is $4$.
While cut-relative $\mathsf{SNDP}$ and path-relative $\mathsf{SNDP}$ are not in general equivalent, they are closely related. Dinitz et al. [9] showed (see Lemma A.1 in [9]) that one can give a cut-based formulation for path-relative $\mathsf{SNDP}$, where $H$ is feasible iff $|\delta_{H}(S)|\geq\min\bigl{\{}f(S),|\delta_{G}(S)|\}$ for every $S\subseteq V$ such that $G[{S}]$ and $G[{V^{\prime}-S}]$ are connected graphs, where $V^{\prime}\subseteq V$ is the vertex-set of the component containing $S$. That is, we require the cut-constraints of $\mathsf{CR}\text{-}\mathsf{SNDP}$ to hold for a suitable collection of node-sets (as opposed to all node-sets in $\mathsf{CR}\text{-}\mathsf{SNDP}$). This immediately implies that $\mathsf{CR}\text{-}\mathsf{SNDP}$ and path-relative $\mathsf{SNDP}$ are equivalent when the underlying graph is a (capacitated) complete graph. Also, as mentioned earlier, path-relative $k\text{-}\mathsf{ECSS}$ and cut-relative $k\text{-}\mathsf{ECSS}$ are equivalent.
Lemma A.1 (Follows from [8]).
Path-relative $k\text{-}\mathsf{ECSS}$ and cut-relative $k\text{-}\mathsf{ECSS}$ are equivalent.
Proof A.2.
This was shown by [8]. We include a proof for completeness. We have already seen that a feasible solution to $\mathsf{CR}\text{-}\mathsf{SNDP}$ is also feasible for path-relative $\mathsf{SNDP}$. Now suppose that $H\subseteq E$ is a feasible solution to path-relative $k\text{-}\mathsf{ECSS}$. Suppose $\delta_{H}(S)<\min\bigl{\{}k,|\delta_{G}(S)|\bigr{\}}$ for some set $\emptyset\neq S\subsetneq V$. Consider $F=|\delta_{H}(S)|$, and let $e=uv\in\delta_{G-F}(S)$. Then, $|F|<k$ and $G-F$ has a $u$-$v$ path, but $H-F$ has no $u$-$v$ path, which contradicts that $H$ is feasible for path-relative $k\text{-}\mathsf{ECSS}$.
Appendix B Proof of part (b) of Theorem 4.7
We mimic the proof of part (a). Recall that $Z=\delta_{G}(A)$, $(f_{1},G_{1},F_{1})=\bigl{(}f_{A},G[{A}],F(A)\bigr{)}$ and $(f_{2},G_{2},F_{2})=\bigl{(}f_{\overline{A}},G[{\overline{A}}],F(\overline{A})\bigr{)}$.
We extend the notion of deficiency to fractional solutions, or equivalently weighted graphs. Given graph $D=(V_{D},E_{D})$, weights $w\in\mathbb{R}_{+}^{E_{D}}$ and cut-requirement function $h:2^{V_{D}}\mapsto\mathbb{Z}$, define the weighted-deficiency of a set $S\subseteq V$ with respect to $(h,w)$ to be $\operatorname{wdef}_{h,w}(S):=h(S)-w\bigl{(}\delta_{D}(S)\bigr{)}$. If $w_{e}=1$ for all $e\in Z$, then as in part (b) of Lemma 4.5, we have $\operatorname{wdef}_{f_{1},w^{A}}(T)=\max\bigl{\{}\operatorname{wdef}_{f,w}(T),\operatorname{wdef}_{f,w}(T\cup\overline{A})\bigr{\}}$ for $T\subseteq A$. Similarly, by considering $f_{2}$, the restriction of $f$ to $\overline{A}$, we have $\operatorname{wdef}_{f_{2},w^{\overline{A}}}(T)=\max\bigl{\{}\operatorname{wdef}_{f,w}(T),\operatorname{wdef}_{f,w}(T\cup A)\bigr{\}}$ for $T\subseteq\overline{A}$.
Let $x\in\mathbb{R}_{+}^{E}$. Suppose first that $x$ is feasible to $(\text{{CRLP}}_{f,G})$. Since $A$ is a small-$(f,G)$-cut, we must have $x_{e}=1$ for all $e\in Z$. Consider any $\emptyset\neq T\subsetneq A$. We have $x\bigl{(}\delta_{G}(T)\bigr{)}\geq\min\bigl{\{}f(T),|\delta_{G}(T)|\bigr{\}}$, so since $x_{e}=1$ for all $e\in Z$, we have $x^{A}\bigl{(}\delta_{G_{1}}(T)\bigr{)}\geq\min\bigl{\{}f(T)-|\delta_{Z}(T)|,|\delta_{G_{1}}(T)|\bigr{\}}$. We also have $x\bigl{(}\delta_{G}(T\cup\overline{A})\bigr{)}\geq\min\bigl{\{}f(T\cup\overline{A}),|\delta_{G}(T\cup\overline{A})|\bigr{\}}$, and so $x^{A}\bigl{(}\delta_{G_{1}}(T)\bigr{)}\geq\min\bigl{\{}f(T\cup\overline{A})-|\delta_{Z}(T\cup\overline{A})|,|\delta_{G_{1}}(T)|\bigr{\}}$. Combining these, we obtain that $x^{A}\bigl{(}\delta_{G_{1}}(T)\bigr{)}\geq\min\bigl{\{}f_{1}(T),|\delta_{G_{1}}(T)|\bigr{\}}$. Thus shows that $x^{A}$ is feasible to $(\text{{CRLP}}_{(f_{1},G_{1})})$.
Again, a symmetric argument interchanging the roles of $A,f_{1},F_{1},G_{1}$ with $\overline{A},f_{2},F_{2},G_{2}$ shows that $x^{\overline{A}}$ is a feasible solution to $(\text{{CRLP}}_{(f_{2},G_{2})})$.
Now suppose that $x_{e}=1$ for all $e\in Z$, $x^{A}$ is feasible to $(\text{{CRLP}}_{(f_{1},G_{1})})$, and $x^{\overline{A}}$ is feasible to $(\text{{CRLP}}_{(f_{2},G_{2})})$. Consider any $\emptyset\neq T\subsetneq V$. Suppose $T\subseteq A$. If $T$ is a small-$(f_{1},G_{1})$-cut, then $x_{e}=1$ for all $e\in\delta_{G_{1}}(T)$; since $x_{e}=1$ for all $e\in Z$, this implies that $x_{e}=1$ for all $e\in\delta_{G}(T)$. Otherwise, we have $\operatorname{wdef}_{f_{1},x^{A}}(T)\leq 0$, since $x^{A}$ is feasible to $(\text{{CRLP}}_{(f_{1},G_{1})})$, and so $\operatorname{wdef}_{f,x}(T)\leq\operatorname{wdef}_{f_{1},x^{A}}(T)\leq 0$. This shows that the constraints for sets $T\subseteq A$ in $(\text{{CRLP}}_{f,G})$ is satisfied. A symmetric argument shows that the same holds for sets $T\subseteq\overline{A}$.
Now consider $T$ such that $T\cap A$ and $T\cap\overline{A}$ are both non-empty. As in the proof of part (a), since $\operatorname{wdef}_{f,x}$ is weakly supermodular and symmetric, we can find sets $X\subseteq A$, $Y\subseteq\overline{A}$ such that $\operatorname{wdef}_{f,x}(T)+\operatorname{wdef}_{f,X}(A)\leq\operatorname{wdef}_{f,x}(X)+\operatorname{wdef}_{f,x}(Y)$ and $\delta_{G}(T)\subseteq\delta_{G}(X)\cup\delta_{G}(Y)\cup Z$. We showed above that the constraints for $X$ and $Y$ in $(\text{{CRLP}}_{f,G})$ are satisfied, which can be stated equivalently as $\operatorname{wdef}_{f,x}(S)\leq\max\bigl{\{}0,\operatorname{def}_{f,G}(S)\bigr{\}}$ for $S\in\{X,Y\}$. Also, $\operatorname{wdef}_{f,x}(A)=\operatorname{def}_{f,G}(A)>0$. So for $S\in\{X,Y\}$, we have
$\operatorname{wdef}_{f,x}(S)\leq\max\bigl{\{}0,\operatorname{def}_{f,G}(S)\bigr{\}}\leq\max\bigl{\{}0,\operatorname{def}_{f,G}(A)\bigr{\}}=\operatorname{def}_{f,G}(A)=\operatorname{wdef}_{f,x}(A)$ (9)
where the second inequality is because $A$ is a maximum-deficiency set. If both $X$ and $Y$ are small-$(f,G)$-cuts, then we have $x_{e}=1$ for all $e\in\delta_{G}(X)\cup\delta_{G}(Y)\cup Z$, and hence $x_{e}=1$ for all $e\in\delta_{G}(T)$. Otherwise, $\operatorname{wdef}_{f,x}(T)+\operatorname{wdef}_{f,X}(A)\leq\operatorname{wdef}_{f,x}(X)+\operatorname{wdef}_{f,x}(Y)$, coupled with (9) shows that $\operatorname{wdef}_{f,x}(T)\leq\min\bigl{\{}\operatorname{wdef}_{f,x}(X),\operatorname{wdef}_{f,x}(Y)\bigr{\}}$, which is at most 00 since at least one of $X$, $Y$ is not a small-$(f,G)$-cut. This shows that $x$ is feasible to $(\text{{CRLP}}_{(f,G)})$. ∎